### v2

# Tree-structured Ising models under mean parameterization

Abstract
We assess advantages of expressing tree-structured Ising models via their mean parameterization rather than their commonly chosen canonical parameterization. This includes fixedness of marginal distributions, often convenient for dependence modeling, and the dispelling of the intractable normalizing constant otherwise hindering Ising models. We derive an analytic expression for the joint probability generating function of mean-parameterized tree-structured Ising models, conferring efficient computation methods for the distribution of the sum of its constituent random variables. The mean parameterization also allows for a stochastic representation of Ising models, providing straightforward sampling methods. We furthermore show that Markov random fields with fixed Poisson marginal distributions may act as an efficient and accurate approximation for tree-structured Ising models, in the spirit of Poisson approximation.
Keywords: Multivariate Bernoulli distribution, dependence tree, undirected graphical models, binary Markov random fields, Poisson approximation, Jungle model, Autologistic model, Markov-Bernoulli distribution.
1 Introduction
The Ising model stems from work in physics, first introduced in [Ising, 1924], where it was proposed as a model to describe macroscopic ferromagnetism phenomena from elementary inter-atomic electromagnetic interaction; for historical remarks on the Ising model, see [Ising et al., 2017] and [Budrikis, 2024]. For an introduction to the Ising model in physics, see [Okounkov, 2022]. Nowadays, however, the Ising model has overreached beyond physics applications and become a most notorious graphical model with multidisciplinary applications. Textbooks on probabilistic graphical models (among which [Koller and Friedman, 2009], [Maathuis et al., 2018], [Theodoridis, 2015] and [Wainwright and Jordan, 2008]) invariably present the model as a classic example of Markov random fields (MRF). Due to its multidisciplinary reach, the Ising model has been referred to under other names, varying upon some branches of the literature. Research articles stemming from the work in [Besag, 1974] will call it the autologistic (or auto-logistic) model; articles in statistical science may call it the binary MRF, for instance [Kaiser et al., 2014] and [Venema, 1992]; and has been employed under the name Jungle model in credit risk modeling, starting from [Molins and Vives, 2016]. Although some of the models have slight variations in their formulation, they all refer to the same multivariate distribution. This will be made evident by our first main result, Theorem 1, in Section 2. Throughout, we let the name Ising model subsume all of these models, given its ubiquity in the literature and the anteriority of [Ising, 1924].
Originating from physics, the Ising model is almost always employed in its formulation as a member of the exponential family, since the parameterization coming along then has a physical interpretation. For example, in the most elementary one-parameter Ising model, this parameter represents the inverse temperature of the ferromagnetic metal ([Kindermann and Snell, 1980]). However, in most other contexts, these parameters lack direct interpretability in a distributional sense. In this vein, [Caragea and Kaiser, 2009] proposed an alternate parameterization, still in exponential form, rendering pseudo-marginal parameters; their accuracy is nonetheless influenced by introducing dependence. Another drawback, affecting any MRF in its exponential formulation, is the presence of an intractable normalizing constant.
In a dependence modeling context, two crucial distributional ingredients are the marginal distributions and the pairwise correlations. Another desirable feature is avoiding the magnitude of the individually observed phenomena changing with dependence. The parameterization inherited from physics does not provide these ingredients nor this feature as varying one parameter of the distribution influences all marginal distributions and pairwise correlations, unless one makes the restrictive assumption of "no external field". This assumption bears limited significance in fields of applications other than physics and imposes symmetric marginal distributions, limiting to a dependence modeler. In its exponential form, the Ising model arises as an MRF conceived through the fixed-conditional paradigm, as fully discussed in [Besag, 1974]; and this what prevents such properties.
In Chapter 3 of [Wainwright and Jordan, 2008], the authors examine an inversion of that paradigm, letting marginal distributions be fixed, yielding another parameterization referred to as the mean parameterization. The paradigm inversion had also been discussed and performed for a lattice MRF in [Pickard, 1980]; this idea also underlied the design of the MRF in [Pickard, 1977]. In [Côté et al., 2025], a MRF with Poisson marginal distributions is introduced under the fixed-marginal paradigm. The particular case of the Ising model is treated in Example 3.8 of [Wainwright and Jordan, 2008]. Under this new paradigm, parameters correspond to the distributional ingredients most suitable for dependence modeling and the separation feature is displayed. The intractable normalizing constant is also dispelled. Our search for further findings on the Ising model under this paradigm has however provided only a handful of research works, for example [Dobra, 2003] for general MRFs in the context of contingency tables. This paper puts forth new results obtained due to the use of this parameterization.
A first main objective of this paper is to discuss the different parameterizations of the Ising model and how they are intertwined. This allows a better understanding of the assumptions underlying the model and highlights the advantages of each one. An example is the mean parameterization shedding light on the notorious Markov-Bernoulli, introduced in [Edwards, 1960] and notably studied in [Gani, 1982], as a special case of an Ising model. To have an outlook on the Ising model from different perspectives uncovers and connects many of its distributional properties. Some ways to envision the model may elude to a researcher fixated on one parameterization; our results in Section 5 and 6 have value beyond their computational aspect in that sense.
A second main objective of this paper is to put forth the nice properties of Ising models when defined on trees, especially as they allow for a mean parameterization with simple constraints. Trees also offer sequential conditioning, as highlighted in [Côté et al., 2025]. They are most useful in a dependence modeling context as there exist efficient tree-structured learning algorithms [Chow and Liu, 1968], see notably [Bresler and Karzand, 2020] and [Nikolakakis, 2021] for a deep investigation. We provide theoretical results of interest allowing a deeper comprehension of the dynamics at play in tree-structured Ising models and their relevance for computational purposes.
The paper is constructed as follows. In Section 2, we present and connect the different parameterizations of Ising models and discuss the advantages and drawbacks of each. We derive the joint probability generating function (pgf) of a tree-structured Ising model, a novel result, in Section 3 and present efficient computation methods leveraging this result in Section 4. In Section 5, we discuss how the mean parameterization and the tree-structure combined provide a stochastic representation and a straightforward sampling method. In Section 6, we examine how the family of MRFs introduced in [Côté et al., 2025] may approximate tree-structured Ising models. Section 7 presents a numerical example.
2 Parameterizations
Let $d\in\mathbb{N}^{*}=\mathbb{N}\backslash\{0\}$ and $\mathcal{V}=\{1,\ldots,d\}$. Define $\mathscr{V}$ as the power set of $\mathcal{V}$, excluding the empty set $\varnothing$. For example, if $\mathcal{V}=\{1,2,3\}$, then $\mathscr{V}=\{\{1\},\{2\},\{3\},\{1,2\},\{1,3\},\{2,3\},\{1,2,3\}\}$. Any $d$-dimensional vector of random variables $\bm{J}=(J_{v},\,v\in\mathcal{V})$ with each of its components following a Bernoulli distribution has a joint probability mass function (pmf) expressible as
$p_{\bm{J}}(\bm{x})=\prod_{W\in\mathscr{V}\cup\{\varnothing\}}\mathrm{Pr}\left(\bigcap_{v\in\mathcal{V}}\left\{J_{v}=\mathbbm{1}_{\{v\in W\}}\right\}\right)^{\left(\prod_{v\in W}x_{v}\right)\left(\prod_{u\in\mathcal{V}\backslash W}(1-x_{u})\right)},\quad\bm{x}\in\{0,1\}^{d}.$ (1)
The pmf in (1) describes the multivariate Bernoulli distribution, a family which encompasses all distributions with Bernoulli marginals. [^1] One may assume that such a feature would naturally follow given the generality of the distribution’s name, by necessity for consistent terminology. This is not always the case, although: for example, the multivariate Poisson distribution does not encompass all joint distributions with Poisson marginals, see e.g. [Inouye et al., 2017]. For further insight on the multivariate Bernoulli distribution, see [Dai et al., 2013]. After manipulations, given some equivalences [^2] Lemma 3.1 of [Dai et al., 2013] provides these equivalences; we restate them in Appendix B for the reader’s convenience. between the probabilities $\mathrm{Pr}(\bigcap_{v\in W}\{J_{v}=1\},\bigcap_{v\in\mathcal{V}\backslash W}\{J_{v}=0\})$ and a vector of parameters $\bm{\eta}=(\eta_{W},\,W\in\mathscr{V})\in\mathbb{R}^{2^{d}-1}$, the joint pmf in (1) is rewritten as a member of the exponential family as
$p_{\bm{J}}(\bm{x})=\mathrm{exp}\left({\sum_{W\in\mathscr{V}}\eta_{W}\prod_{v\in W}x_{v}-A(\bm{\eta})}\right),\quad\bm{x}\in\{0,1\}^{d},$ (2)
where $A(\bm{\eta})$ serves as a normalizing constant. Therefore, (2) also describes all possible joint pmfs for a vector of Bernoulli random variables. Following [Dai et al., 2013]’s terminology, the vector of parameters $\bm{\eta}$ is called the natural parameters, making (2) the natural representation of the joint pmf of $\bm{J}$. Natural parameters of the $i$th order are the ones associated to elements of $\mathscr{V}$ which themselves comprise $i$ elements, $i\in\{1,\ldots,d\}$.
Ising models generate a subclass of multivariate Bernoulli distributions. In the upcoming subsections, we present three formulations of this model: natural, canonical, and – when tree structured – mean parameterized. Some notation surrounding graphs and trees are required beforehand.
[…]
Labelling one vertex, say $r$, as the root gives the rooted tree $\mathcal{T}_{r}$. When a tree is rooted, the following filiation relations between vertices become well defined, according to that root: the descendants of $u$, $u\in\mathcal{V}$, are the set of vertices such that $u$ is on their path to the root, $\mathrm{dsc}(u)=\{j\in\mathcal{V}:(v,w)\in\mathrm{path}(r,j),\,w\in\mathcal{V}\}$; the children of $u$ are the set of vertices that are both its descendant and its neighbor, $\mathrm{ch}(u)=\{j\in\mathcal{V}:j\in\mathrm{nei}(u)\cap\mathrm{dsc}(u)\}$; the parent of $u$ is its sole neighboring vertex that is not its children, $\mathrm{pa}(u)\in\mathrm{nei}(u)\backslash\mathrm{ch}(u)$. The root has no parent since all other vertices are its descendants. One may refer to Chapter 3.3 of [Saoub, 2021] for further insight on rooted trees.
2.1 Natural parameterization
The Ising model is a special case of the multivariate Bernoulli distribution where higher-order interactions between random variables are dismissed.
Definition 1 (Ising model).
A multivariate Bernoulli distribution corresponds to that of an Ising model if its natural parameters of third order and more are all zero. An Ising model defines itself on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, whence non-zero second order natural parameters dictates its edges, that is, $(u,v)\in\mathcal{E}$ if and only if $\eta_{\{u,v\}}>0$. An Ising model is tree-structured if $\mathcal{G}$ is a tree.
The probability mass function of an Ising model $\bm{J}=(J_{v},\,v\in\mathcal{V})$ defined on graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and with vector of parameters $\bm{\eta}=(\eta_{W}:\;W\in\mathscr{V},|W|\leq 2)\in\mathbb{R}^{2d-1}$ is given by
$p_{\bm{J}}(\bm{x})=\mathrm{exp}\left({\sum\limits_{v\in\mathcal{V}}\eta_{v}x_{v}+\sum\limits_{(u,v)\in\mathcal{E}}\eta_{\{u,v\}}x_{u}x_{v}-A(\bm{\eta})}\right),\quad\bm{x}\in\{0,1\}^{d}.$ (3)
We will henceforth refer to (3) as the joint pmf in its natural formulation, and $\bm{\eta}$ are thus the natural parameters as in (2). The model in (3) is called the auto-logistic model in [Besag, 1974] and the Jungle model in some financial applications ([Emonti and Fontana, 2025], [Molins and Vives, 2016]).
A consequence of zero higher-order natural parameters is conditional independence properties making the Ising model a MRF. Chapter 4.2 of [Cressie and Wikle, 2015] provides the following definition of a MRF.
Definition 2 (MRF).
A vector of random variables $\bm{X}=(X_{v},\,v\in\mathcal{V})$ is a MRF encrypted on graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ if it satisfies the local Markov property on that graph, meaning
$X_{u}\perp\!\!\!\perp X_{v}|\{X_{j}:(u,j)\in\mathcal{E}\}$ (4)
[…]
The Hammersley-Clifford theorem states that a vector of random variables, for which every joint probability is non-zero, is a MRF encrypted on graph $\mathcal{G}$ if and only if its distribution is Gibbs with respect to $\mathcal{G}$.
Definition 3 (Gibbs distribution).
A distribution is Gibbs with respect to graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ if it is characterized by a pdf/pmf of the form
$p(\bm{x})=\frac{1}{z}\prod_{W\in\mathcal{C}}\varphi_{W}((x_{v}:v\in W)),\quad\bm{x}\in\mathbb{R},$ (5)
for some functions $\{\varphi_{W},\,W\in\mathcal{C}\}$, where $\mathcal{C}$ is the set of cliques of the graph and $z$ is a normalizing constant. Cliques are sets of vertices such that every vertex is a neighbor to every other.
Gibbs distributions also satisfy the global Markov property, see Chapter 3 of [Lauritzen, 1996]. On a tree, the global Markov property amounts to the following conditional independencies:
$X_{u}\perp\!\!\!\perp X_{v}|\;X_{j},\quad\text{with vertex $j$ on }\mathrm{path(u,v)},\quad\text{for every }u,v\in\mathcal{V},\,(u,v)\not\in\mathcal{E}.$ (6)
We address in the theorem below the link between Ising models and MRFs with Bernoulli marginal distributions.
Theorem 1.
Under the assumption that joint probabilities are non-zero for every configuration on $\{0,1\}^{d}$, all tree-structured MRFs with Bernoulli marginal distributions are Ising models.
Proof.
This observation is made by bringing together the definition of a Gibbs distribution and that of an Ising model. On a tree, cliques contain one or two vertices only; therefore, for the pmf given in (2) to factorize over the cliques, the associated Bernoulli multivariate distribution must have null natural parameters of third order or more. It would otherwise be algebraically impossible to factorize over the cliques in any fashion. Hence, it corresponds to the pmf of an Ising model, as given in (8). ∎
Note that the statement would not hold for graphs in general, as they may have higher-dimension cliques allowing the presence of higher-order natural parameters meanwhile satisfying Gibbs factorization. Moreover, Theorem 1 should not be wrongly interpreted as: "All tree-structured graphical models with Bernoulli marginals are Ising models". For example, the binary graphical model proposed in the BDgraph package [Mohammadi and Wit, 2019], a model which some researchers have wrongly referred to as a Ising model [Denuit and Robert, 2022], does not satisfy the conditional independence property (4) characterizing a MRF. This model is obtained from a marginal transformation of a Gaussian MRF; this transformation dispels the conditional independence property. It is a member of the family of copula Gaussian graphical models introduced in [Dobra and Lenkoski, 2011].
2.2 Canonical parameterization
[…]
$p_{\bm{J}}(\bm{y})=\mathrm{exp}\left({\sum\limits_{v\in\mathcal{V}}\theta_{v}y_{v}+\sum\limits_{(u,v)\in\mathcal{E}}\theta_{\{u,v\}}y_{u}y_{v}-Z(\bm{\theta})}\right),\quad\bm{y}\in\{-1,1\}^{d}.$ (7)
This parameterization is usually the one preferred in the physics literature, and an uninitiated reader may be oblivious to this subtlety regarding the support when browsing the literature on Ising models. Note that (7) is equivalent to
$p_{\bm{J}}(\bm{x})=\mathrm{exp}\left({\sum\limits_{v\in\mathcal{V}}\theta_{v}\frac{x_{v}-1}{2}+\sum\limits_{(u,v)\in\mathcal{E}}\theta_{\{u,v\}}\frac{x_{u}-1}{2}\frac{x_{v}-1}{2}-Z(\bm{\theta})}\right),\quad\bm{x}\in\{0,1\}^{d}.$ (8)
The Ising model in (7), is generally referred to with the qualifier free boundary in physics-related works to emphasize that no component of $\bm{J}$ are degenerate ([Kindermann and Snell, 1980], Chapter 5). By Theorem 1, there is no ambiguity that the pmf in (8) describes the same distribution as the one in (3), with some equivalence between parameters $\bm{\eta}$ and $\bm{\theta}$.
As we stated in the Introduction, the Ising model has many different names in the literature, and one could think that the terminology may actually follow the support of the distribution, in the same manner that the Rademacher and Bernoulli distributions bear different names although they refer to the same stochastic behavior, only on different supports. For example, binary MRFs most often refer to the distribution on $\{0,1\}^{d}$. This is, however, not exactly consistent in the literature: sources may refer to the distribution with support on $\{0,1\}^{d}$ as Ising models (e.g. [Dai et al., 2013], [Denuit and Robert, 2022]) and other sources refer to the distribution with support on {$,1\}$ 1 } as binary MRFs (e.g. [Venema, 1992]). This is why we have chosen to refer to all as Ising models, given the anteriority of [Ising, 1924] and due to its predominance in the literature. This also emphasizes that any distributional consideration relates to both cases, up to a mapping between parameters $\bm{\eta}$ and $\bm{\theta}$ — as the equivalence between (7) and (8) highlights.
2.3 Centered parameterization
In [Caragea and Kaiser, 2009], the authors suggests transforming the vertex-idiosyncratic parameters from the natural parameterization into a vector $\bm{\kappa}=(\kappa_{v},\,v\in\mathcal{V})\in(0,1)^{d}$, defined so that
$\eta_{v}=\ln\left(\frac{\kappa_{v}}{1-\kappa_{v}}\right)-\sum_{j\in\mathrm{nei}(v)}\kappa_{j}\eta_{\{v,j\}},\quad v\in\mathcal{V}.$
Such is referred to as the centered parameterization of (3), which becomes
$p_{\bm{J}}(\bm{x})=\mathrm{exp}\left(\sum_{v\in\mathcal{V}}x_{v}\ln\left(\frac{\kappa_{v}}{1-\kappa_{v}}\right)+\sum_{(u,v)\in\mathcal{E}}\eta_{\{u,v\}}(x_{u}-\kappa_{u})(x_{v}-\kappa_{v})-B(\bm{\eta},\bm{\kappa})\right),\quad\bm{x}\in\{0,1\}^{d},$ (9)
with $\bm{\kappa}\in(0,1)^{d}$ and $\{\eta_{e}:e\in\mathcal{E}\}\in\mathbb{R}^{d-1}$, and where $B(\bm{\eta},\bm{\kappa})$ is a normalizing constant. [Caragea and Kaiser, 2009] also accounted for covariates, in an autologistic regression perspective, but this is not the focus of the present paper. The advantage of the centered parameterization is that the vector $\bm{\kappa}$ becomes an approximation of the marginal distributions’ means when the interactions given by $\{\eta_{e}:e\in\mathcal{E}\}$ are not too strong. The centered parameterization is further discussed for general MRFs in [Kaiser et al., 2012].
2.4 Mean parameterization
The Ising model is commonly referred to and employed in its canonical parameterization (8), as the parameters then have a physical interpretation in the context of ferromagnetism, from which the model originates [Ising, 1924]. Given its close similarity, the natural parameterization is also often used when modeling dependence between binary random variables, such as in [Denuit and Robert, 2022] and [Lindberg et al., 2016]. The centered parameterization has also been employed, for example in [Kaiser et al., 2014] and [Wakeland, 2017]. However, these parameterizations come with some major drawbacks for dependence modeling purposes. First, the normalization constants $A(\bm{\eta})$, $Z(\bm{\theta})$ and $B(\bm{\eta},\bm{\kappa})$, are arduously computable as they involve a summation over every combination on $\{0,1\}^{d}$. Second, for the natural formulation, the marginal distributions are not fixed: an alteration of the dependence between $u$ and $v$ through $\eta_{(u,v)}$ results in a change of the marginal distributions associated to each vertex in the model. This is also true for the canonical parameterization, albeit for some special cases (discussed below), and the centered parameterization, wherein the goodness of the approximation given by $\bm{\kappa}$ is influenced by $\{\eta_{e},\,e\in\mathcal{E}\}$. It resulted in many algorithms, often either complex or approximative, to compute likelihoods or marginal probabilities; see [Wainwright and Jordan, 2008] and [Theodoridis, 2015] for an exposition.
[…]
$p_{\bm{J}}(\bm{x})=p_{J_{r}}(x_{r})\prod_{v\in\mathcal{V}\backslash\{r\}}p_{J_{v}|J_{\mathrm{pa}(v)}=x_{\mathrm{pa}(v)}}(x_{v})=p_{J_{r}}(x_{r})\prod_{v\in\mathcal{V}\backslash\{r\}}\frac{p_{J_{v},J_{\mathrm{pa}(v)}}(x_{v},x_{\mathrm{pa}(v)})}{p_{J_{\mathrm{pa}(v)}}(x_{\mathrm{pa}(v)})}.$ (10)
From the mean parameterization, marginal pmfs $p_{J_{v}}$, for $v\in\mathcal{V}$, are thus expressed as $p_{J_{v}}(x)=q_{v}^{x_{v}}(1-q_{v})^{1-x_{v}}$, and
$p_{J_{v},J_{\mathrm{pa}(v)}}(x_{v},x_{\mathrm{pa}(v)})=p_{J_{v}}(x_{v})p_{J_{\mathrm{pa}(v)}}(x_{\mathrm{pa}(v)})+\alpha_{(\mathrm{pa}(v),v)}(-1)^{x_{v}+x_{\mathrm{pa}(v)}}\sigma_{(\mathrm{pa}(v),v)}.$ (11)
The joint pmf in (10) becomes
$p_{\bm{J}}(\bm{x})$ $=p_{J_{r}}(x_{r})\prod_{v\in\mathcal{V}\backslash\{r\}}\frac{p_{J_{v}}(x_{v})p_{J_{\mathrm{pa}(v)}}(x_{\mathrm{pa}(v)})+\alpha_{(\mathrm{pa}(v),v)}(-1)^{x_{v}+x_{\mathrm{pa}(v)}}\sigma_{(\mathrm{pa}(v),v)}}{p_{J_{\mathrm{pa}(v)}}(x_{\mathrm{pa}(v)})},$ (12)
where $\sigma_{(u,v)}=\sqrt{q_{u}q_{v}(1-q_{u})(1-q_{v})}$, for $u,v\in\mathcal{V}$, a notational convention kept throughout the paper. Obviously, one may choose any root $r$ to define the filiation relations in (12) and will obtain the same expression for the joint pmf, perhaps only different in appearance. The rooting serves only to specify an order of conditioning.
For every multiplicand in (12) to be strictly positive, one requires (11) to always be strictly positive, meaning
$\alpha_{(\mathrm{pa}(v),v)}\in\left(-\min\left(\sqrt{\tfrac{q_{\mathrm{pa}(v)}q_{v}}{(1-q_{\mathrm{pa}(v)})(1-q_{v})}},\sqrt{\tfrac{(1-q_{\mathrm{pa}(v)})(1-q_{v})}{q_{\mathrm{pa}(v)}q_{v}}}\right),\min\left(\sqrt{\tfrac{(1-q_{\mathrm{pa}(v)})q_{v}}{q_{\mathrm{pa}(v)}(1-q_{v})}},\sqrt{\tfrac{q_{\mathrm{pa}(v)}(1-q_{v})}{(1-q_{\mathrm{pa}(v)})q_{v}}}\right)\right)$ (13)
for every $v\in\mathcal{V}$, thus bounding the possible dependence schemes according to local interactions. Proposition 4.1 of [Wainwright and Jordan, 2008] indicates that the constraints on $\bm{\alpha}$ given by (13), ensuring locally consistent distributions, coincide with those of globally realizable distributions. In other words, the marginal problem on tree $\mathcal{T}$ (see [Jha and Nechita, 2025]) is trivial in the Bernoulli case. These constraints on $\bm{\alpha}$ form a convex hull and are greatly discussed in Chapter 3 of [Wainwright and Jordan, 2008]. Henceforth, we will denote by $\Lambda(\bm{q})$ the set of admissible $\bm{\alpha}$ for given $\bm{q}\in(0,1)^{d}$. Note that we have considered an open interval in (13) to exclude the limit cases where some joint probabilities are null; and hence making sure the Hammersley-Clifford theorem and Theorem 1 apply. One recognizes in (12) a joint pmf belonging to a Gibbs distribution factorizing on $\mathcal{T}$, according to Definition 3. Given the Hammersley-Clifford theorem and Theorem 1, we have that the natural formulation in (3), the canonical formulation in (8) and the centered formulation in (9) are necessarily equivalent expressions to the pmf given in (12), up to some relationship between the $\bm{\eta}$, $\bm{\theta}$ or $\bm{\kappa}$ parameters and the $\bm{q}$ and $\bm{\alpha}$ parameters of the latter.
An advantage of the tree structure is that the constraints on $\bm{\alpha}$ in (13) are actually simple enough to be derived. For graphs, mean parameters would have to satisfy more complex constraints, not only for the clique-based multiplicands in (12) to be positive but also for the higher-order natural parameters that would result to be zero. The latter cause is really what leads to their intractability, due to that of the mapping between both parameterizations (see Appendix C) and because it would vary with the shape of the graph. In Chapter 3.4 of [Wainwright and Jordan, 2008], the authors explain that the problems of marginalization and computation of $A(\bm{\eta})$ for exponential-form MRFs may actually be interpreted as mappings between their natural and mean parameterizations.
The mean parameterization for tree-structured Ising models has the advantage of explicitly stating the marginals and the pairwise correlations. In addition to this enhancement in intelligibility, it is much more convenient in practical terms: changing the dependence between two neighboring components of $\bm{J}$ via the $\bm{\alpha}$ parameters does not affect the marginals. One also notices that there are no embarrassing normalization constant. One may consult Chapter 5 of [Kindermann and Snell, 1980] for a discussion of the marginal pmfs of tree-structured Ising models under their canonical parameterization. in order to highlight the non-fixedness of marginals and abstruse impacts on correlation when employing the natural parameterization. Note that [Teugels, 1990] established an alternate parameterization of the joint pmf of the multivariate Bernoulli distribution (2), directly linked to the marginal distributions and joint moments of $\bm{J}$, and this parameterization is used in [Fontana and Semeraro, 2018]. In Appendix C, we express $\bm{q}$ and $\bm{\alpha}$ as functions of $\bm{\eta}$. Theorem 2.1 of [Filiz et al., 2012] attests the existence of a unique converse mapping.
The mean parameterization highlights overlooked connections between the Ising model and other models proposed in the literature. For instance, we notice that the Markov Bernoulli distribution, introduced in [Edwards, 1960] and studied in [Gani, 1982], corresponds to an Ising model on a series tree. Their equivalence is clear from Theorem 1. In a similar vein, we remark that the tree graphical model with conditional independence studied in [Padmanabhan and Natarajan, 2021] is an Ising model, again under its mean parameterization.
As explained in [Côté et al., 2025], there are two paradigms to define MRFs. On one hand, one may specify the conditional distributions when every neighbor is known; or on the other hand, fix the marginal distributions. When conditional distributions are specified, marginals are usually hardly tractable; and the converse holds for fixed-marginals. Hitherto, MRFs with fixed-conditional distributions have been extensively studied, such as the auto-models of [Besag, 1974]. Both paradigms do not usually lead to the same MRFs. For example, the auto-Poisson MRF in [Besag, 1974], defined through conditional distributions, differs from the MRF obtained by fixing Poisson marginals in [Côté et al., 2025]. Values taken by the neighbor’s random variables within the tree in Besag’s auto-Poisson model affect the marginal distributions: a modification of dependence parameters may impact significantly the marginal distributions. The Ising model constitutes an exception to this discrepancy: both the fixed-conditionals and the fixed-marginals approaches lead to the Ising model. The only difference is in the most suitable parameterization ensuing. The natural parameterization explicitly states the conditional distribution, see equation (4.8) in [Besag, 1974]; for fixed-marginals, as we discussed above, mean parameterization is most relevant. Not only that correlation between components of $\bm{J}$ whose underlying vertices are connected by an edge are directly specified, correlations between any pair of components follow straightforwardly, as shown in the next proposition.
Proposition 1.
For a tree-structured Ising model $\bm{J}=(J_{v},\,v\in\mathcal{V})$ defined on tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$, formulated through its mean parameterization with vectors $\bm{q}=(q_{v},\,v\in\mathcal{V})\in(0,1)^{d}$ and $\bm{\alpha}=(\alpha_{e},\,e\in\mathcal{E})\in\Lambda(\bm{q})$, the Pearson correlation coefficient between any two components of $\bm{J}$ is
$\mathrm{Corr}(J_{u},J_{v})=\prod_{e\in\mathrm{path}(u,v)}\alpha_{e},\quad u,w\in\mathcal{V}.$
Proof.
The proof is provided in Appendix A.1. ∎
A common assumption when employing the canonical parameterization is that of "no external field", for which $\theta_{v}=0$ for all $v\in\mathcal{V}$. In this special case, the mapping between canonical and mean parameters is direct: $q_{v}=1/2$ for all $v\in\mathcal{V}$ and $\alpha_{e}=\tanh\theta_{e}$ for all $e\in\mathcal{E}$. For this reason, the ensuing model is also sometimes referred to as the symmetric Ising model. This particular case’s mapping is a well-known fact on the Ising model and often used without proof. One may consult [Nikolakakis et al., 2021], Lemmas 12 and 14, for a complete proof. A mapping between canonical and mean parameters becomes highly arduous as soon as the no-external-field assumption is removed, and one must employ marginalization algorithms to proceed, see [Wainwright and Jordan, 2008] and [Theodoridis, 2015]. The no-external-field assumption is very limiting for a dependence modeler: for practical applications, marginal distributions are rarely exactly symmetric. In the statistical literature on the Ising model, however, this assumption is often made – and the authors hence either assume explicitly or implicitly symmetric marginal distributions. Example of such research articles include [Ravikumar et al., 2010], [Daskalakis et al., 2019], [Bresler and Karzand, 2020] and [Nikolakakis et al., 2021].
3 Joint probability generating function
The joint pgf of a vector of discrete random variables $\bm{X}=(X_{i},i\in\{1,\ldots,d\})$ is given by
$\mathcal{P}_{\bm{X}}(\bm{t})=\mathrm{E}\left[t_{1}^{X_{1}}\cdots t_{d}^{X_{d}}\right],\quad\quad\bm{t}=(t_{1},\ldots,t_{d})\in[-1,1]^{d}.$ (14)
The joint pgf is specific to a multivariate distribution and often allows for a deeper understanding of the stochastic dynamics underlying it. Beyond that, its usefulness for computational purposes will be discussed in Section 4 in two different contexts.
Working with the mean parameterization of the Ising model allows to derive the exact expression of its pgf, which we provide in Theorem 2 below. This is because the mean parameterization allows for direct expressions of marginal and bivariate conditional probabilities, without recourse to marginalization algorithms. Also, the cumbersome normalization constants $A(\bm{\eta})$ and $Z(\bm{\theta})$ of the natural and canonical parameterizations also prevents the derivation of an expression of the joint pgf directly from those.
Theorem 2 (Joint pgf).
Consider $\bm{J}=(J_{v},\,v\in\mathcal{V})$, a tree-structured Ising model, mean-parameterized with vectors of parameters $\bm{q}=(q_{v},\,v\in\mathcal{V})\in(0,1)^{d}$ and $\bm{\alpha}=(\alpha_{e},\,e\in\mathcal{E})\in\Lambda(\bm{q})$, defined on tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$. Choose a vertex $r\in\mathcal{V}$ to act as the root of the tree and let $\sigma_{(u,v)}=\sqrt{q_{u}q_{v}(1-q_{u})(1-q_{v})}$ for any $u,v\in\mathcal{V}$. The joint pgf of $\bm{J}$ is given by
$\mathcal{P}_{\bm{J}}(\bm{t})=(1-q_{r})\prod_{i\in\mathrm{ch}(r)}\zeta_{i}^{\mathcal{T}_{r}}(\bm{t}_{i\mathrm{dsc}(i)})+q_{r}t_{r}\prod_{i\in\mathrm{ch}(r)}\xi_{i}^{\mathcal{T}_{r}}(\bm{t}_{i\mathrm{dsc}(i)}),\quad\bm{t}\in[-1,1]^{d},$
with $\zeta_{v}^{\mathcal{T}_{r}}(\bm{t}_{v\mathrm{dsc}(v)})$ and $\xi_{v}^{\mathcal{T}_{r}}(\bm{t}_{v\mathrm{dsc}(v)})$ recursively defined for every $v\in\mathcal{V}\backslash\{r\}$:
$\zeta_{v}^{\mathcal{T}_{r}}\left(\bm{t}_{v\mathrm{dsc}(v)}\right)$ $=\left(1-q_{v}+\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right)\prod_{i\in\mathrm{ch}(v)}\zeta_{i}^{\mathcal{T}_{r}}\left(\bm{t}_{i\mathrm{dsc}(i)}\right)$
$\quad\quad\quad\quad+\left(q_{v}-\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right)t_{v}\prod_{i\in\mathrm{ch}(v)}\xi_{i}^{\mathcal{T}_{r}}\left(\bm{t}_{i\mathrm{dsc}(i)}\right);$ (15)
$\xi_{v}^{\mathcal{T}_{r}}\left(\bm{t}_{v\mathrm{dsc}(v)}\right)$ $=\left(1-q_{v}-\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right)\prod_{i\in\mathrm{ch}(v)}\zeta_{i}^{\mathcal{T}_{r}}\left(\bm{t}_{i\mathrm{dsc}(i)}\right)$
$\quad\quad\quad\quad+\left(q_{v}+\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right)t_{v}\prod_{i\in\mathrm{ch}(v)}\xi_{i}^{\mathcal{T}_{r}}\left(\bm{t}_{i\mathrm{dsc}(i)}\right),$ (16)
where products taken over an empty set are equal to 111 by convention.
Proof.
Reorganizing the factors in the joint pgf gives
$\mathcal{P}_{\bm{J}}(\bm{t})=\mathrm{E}\left[\prod_{v\in\mathcal{V}}t_{v}^{J_{v}}\right]=\mathrm{E}\left[t_{r}^{J_{r}}\prod_{i\in\mathrm{dsc}(r)}t_{i}^{J_{i}}\right]=\mathrm{E}\left[t_{r}^{J_{r}}\prod_{i\in\mathrm{ch}(r)}\left(t_{i}^{J_{i}}\prod_{j\in\mathrm{dsc}(i)}t_{j}^{J_{j}}\right)\right],$
for $\bm{t}=[-1,1]^{d}$. We condition on $J_{r}$, following a Bernoulli distribution of parameter $q_{r}$ according to mean parameterization, and obtain
$\mathcal{P}_{\bm{J}}(\bm{t})=(1-q_{r})\mathrm{E}\left[\left.\prod_{i\in\mathrm{ch}(r)}\left(t_{i}^{J_{i}}\prod_{j\in\mathrm{dsc}(i)}t_{j}^{J_{j}}\right)\right|J_{r}=0\right]+q_{r}t_{r}\mathrm{E}\left[\left.\prod_{i\in\mathrm{ch}(r)}\left(t_{i}^{J_{i}}\prod_{j\in\mathrm{dsc}(i)}t_{j}^{J_{j}}\right)\right|J_{r}=1\right].$
The global Markov property yields the conditional independencies allowing distributivity of expectations over multiplicands:
$\mathcal{P}_{\bm{J}}(\bm{t})=(1-q_{r})\prod_{i\in\mathrm{ch}(r)}\mathrm{E}\left[\left.t_{i}^{J_{i}}\prod_{j\in\mathrm{dsc}(i)}t_{j}^{J_{j}}\right|J_{r}=0\right]+q_{r}t_{r}\prod_{i\in\mathrm{ch}(r)}\mathrm{E}\left[\left.t_{i}^{J_{i}}\prod_{j\in\mathrm{dsc}(i)}t_{j}^{J_{j}}\right|J_{r}=1\right],$ (17)
for $\bm{t}=[-1,1]^{d}$. Next, we find recursive relations to resolve the conditional expectations in (17). Let us observe that, since $(J_{\mathrm{pa}(v)},J_{v})$ follows a bivariate Bernoulli distribution with dependence parameter $\alpha_{(\mathrm{pa}(v),v)}$ provided the mean parameterization, for every $v\in\mathcal{V}\backslash\{r\}$, the following conditional probabilities may be expressed as such:
$\mathrm{Pr}(J_{v}=0|J_{\mathrm{pa}(v)}=0)=\left(1-q_{v}+\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right);\;\mathrm{Pr}(J_{v}=1|J_{\mathrm{pa}(v)}=0)=\left(q_{v}-\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right);$
$\mathrm{Pr}(J_{v}=0|J_{\mathrm{pa}(v)}=1)=\left(1-q_{v}-\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right);\;\mathrm{Pr}(J_{v}=1|J_{\mathrm{pa}(v)}=1)=\left(q_{v}+\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right).$
The following relations are then provided by conditioning on $J_{v}$ and evoking the global Markov property again:
$\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=0\right]$
$\quad=\left(1-q_{v}+\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right)\prod_{i\in\mathrm{ch}(v)}\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=0\right]$
$\quad\quad+\left(q_{v}-\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right)t_{v}\prod_{i\in\mathrm{ch}(v)}\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=1\right];$ (18)
and
$\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=1\right]$
$\quad=\left(1-q_{v}-\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right)\prod_{i\in\mathrm{ch}(v)}\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=0\right]$
$\quad\quad+\left(q_{v}+\frac{\alpha_{(\mathrm{pa}(v),v)}\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right)t_{v}\prod_{i\in\mathrm{ch}(v)}\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=1\right],$ (19)
with $v\in\mathcal{V}\backslash\{r\}$, $\bm{t}\in[-1,1]$. Comparing (18) and (19) with (15) and (16), noting how both pairs of relations recursively intertwine in the same manner, we have $\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=0\right]=\zeta_{v}^{\mathcal{T}_{r}}(\bm{t}_{v\mathrm{dsc}(v)})$ and $\mathrm{E}\left[\left.t_{v}^{J_{v}}\prod_{j\in\mathrm{dsc}(v)}t_{j}^{J_{j}}\right|J_{\mathrm{pa}(v)}=1\right]=\xi_{v}^{\mathcal{T}_{r}}(\bm{t}_{v\mathrm{dsc}(v)})$. Rewriting the joint pgf in (17) accordingly grants the desired result. ∎
Just as for (12), one may choose any root $r$ according to which the joint pgf in Theorem 2 recursively declines itself and obtain the same end result. To specify a root simply allows for the filiation relations, dictating the order of conditioning in the proof, to be clear and well-defined.
One may numerically implement the joint pgf from Theorem 2 by following Algorithm 1 below. For such matters, it is most convenient to employ weighted adjacency matrices to provide essential information about the underlying tree in a digestible format to the algorithm. For a tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$ and vector of parameters $\bm{\alpha}=(\alpha_{e},\,e\in\mathcal{E})$, the corresponding weighted adjacency matrix is given by
$\bm{A}=(A_{ij},\;i\times j\in\mathcal{V}\times\mathcal{V}),\quad A_{ij}=\left\{\begin{array}[]{ll}\alpha_{(i,j)},&(i,j)\in\mathcal{E};\\ 1,&i=j;\\ 0,&\text{elsewhere}.\end{array}\right.$
Here and for henceforth algorithms, we always assume that weighted adjacency matrices are constructed in topological order: according to a given root, a vertex’s associated row and column index comes after its parent’s. The root has index 1. This allows for the algorithm to easily retrieve the filiation relations hence defined, as the first non-zero element of each vertex’s associated column has row index being their parent’s (except for the root’s).
Input: $\bm{t}=(t)_{k\in\mathcal{V}}$; probability parameters $\bm{q}=(q_{k})_{k\in[d]}$; topologically ordered weighted adjacency matrix $\bm{A}=(A_{ij})_{i\times j\in\mathcal{V}\times\mathcal{V}}$.
Output: PGF of $\bm{J}$ evaluated at $\bm{t}$.
Construct $d\times d$ matrices $\bm{p}^{00}$, $\bm{p}^{01}$, $\bm{p}^{10}$ and $\bm{p}^{11}$ as such :
$p^{00}_{ij}=\left((1-q_{i})(1-q_{j})+A_{ij}\sigma_{(i,j)}\right)\mathbbm{1}_{\{A_{ij}>0\}};\quad$ $p^{01}_{ij}=\left((1-q_{i})q_{j}-A_{ij}\sigma_{(i,j)}\right)\mathbbm{1}_{\{A_{ij}>0\}};$
$p^{10}_{ij}=\left(q_{i}(1-q_{j})-A_{ij}\sigma_{(i,j)}\right)\mathbbm{1}_{\{A_{ij}>0\}};\quad$ $p^{11}_{ij}=\left(q_{i}q_{j}+A_{ij}\sigma_{(i,j)}\right)\mathbbm{1}_{\{A_{ij}>0\}}.$
Let $\bm{\zeta}$ and $\bm{\xi}$ be $d\times d$ matrices of zeros.
for $j=d,\ldots,2$ do
 for $i=1,\ldots,j-1$ do
      Replace $\zeta_{ij}$ and $\xi_{ij}$ by
$\zeta_{ij}$ $=\left((1-q_{j})\prod_{k>j}(\zeta_{kj}+\mathbbm{1}_{\{A_{kj}=0\}})\right)\mathbbm{1}_{\{A_{ij}>0\}},\quad\xi_{ij}=\left(t_{j}q_{j}\prod_{k>j}(\xi_{kj}+\mathbbm{1}_{\{A_{kj}=0\}})\right)\mathbbm{1}_{\{A_{ij}>0\}}.$
      Replace $\zeta_{ji}$ and $\xi_{ji}$ by
$\zeta_{ji}$ $=\frac{p^{00}_{i}j}{(1-q_{i})(1-q_{j})}\zeta_{ij}+\frac{p^{01}_{ij}}{(1-q_{i})q_{j}}\xi_{ij},\quad\xi_{ji}=\frac{p^{10}_{ij}}{q_{i}(1-q_{j})}\zeta_{ij}+\frac{p^{11}_{ij}}{q_{i}q_{j}}\xi_{ij}.$
Compute $\zeta^{(1)}=(1-q_{1})\prod_{k>1}(\zeta_{k1}+\mathbbm{1}_{\{A_{k1}=0\}})$.
Compute $\xi^{(1)}=(t_{1}q_{1}\prod_{k>1}(\xi_{k1}+\mathbbm{1}_{\{A_{k1}=0\}})$.
Compute $\mathcal{P}_{\bm{J}}=\zeta^{(1)}+\xi^{(1)}$.
Return $\mathcal{P}_{\bm{J}}$.
Algorithm 1 Computing the PGF of $\bm{J}$.
4 Pgf-related computation methods
In this section, we showcase the usefulness in regard of computational methods of having a closed-form expression of the joint pgf. We are able to, most efficiently, compute the distribution of the sum $K=\sum_{v\in\mathcal{V}}J_{v}$ and quantities $\mathrm{E}[J_{v}\mathbbm{1}_{\{K=k\}}]$ with $v\in\mathcal{V}$, $k\in\mathbb{N}$, problems which have been studied by previous researchers, and for which their solution is either limited to specific cases or more complex to implement.
4.1 Distribution of the sum
[…]
The joint pgf of $\bm{J}$ allows to derive the pgf of $K$ through the relation
$\mathcal{P}_{K}(t)=\mathcal{P}_{\bm{J}}(t,\ldots,t),\quad t\in[-1,1].$
Hence, Theorem 2 also provides a recursive expression for the pgf of $K$. Specifically for a Markov Bernoulli random vector — which corresponds to an Ising model defined on a series-tree structure, as stated above in Section 2.4 —, the pgf of $K$ is also provided in matrix form in [Edwards, 1960]. The joint pgf given in Theorem 2 allows to circumvent the dimensional issue, while allowing for any tree shape and selection of parameters, through Algorithm 2. It relies on the efficiency of the FFT algorithm, see [Embrechts and Frei, 2009] for more details on this approach.
Input: Vector of marginal parameters $\bm{q}=(q_{v})_{v\in\mathcal{V}}$, topologically ordered weighted adjacency matrix $\bm{A}=(A_{ij})_{i\times j\in\mathcal{V}\times\mathcal{V}}$.
Output: Vector $\bm{p}^{(K)}=(p^{(K)}_{k})_{k\in\{0,\ldots,k_{\max}\}}$ such that $p^{(K)}_{k}$ = $p_{K}(k)$.
Set $n_{\mathrm{FFT}}$ to be a large power of 2 (e.g. $2^{13}$). This determines $k_{\max}$.
Set $\bm{b}=(b_{i})_{i\in\{0,\ldots,n_{\mathrm{FFT}}-1\}}=(0,1,0,0,\ldots,0)$.
Use FFT to compute the discrete Fourier transform $\bm{\phi}^{(b)}$ of $\bm{b}$.
for $\ell=0,1,\dots,n_{\rm FFT}-1$ do
   Compute $\phi_{\ell}^{(K)}=\mathcal{P}_{J}(\phi_{\ell}^{(b)},\ldots,\phi_{\ell}^{(b)})$ by employing Algorithm 1.
Use fft to compute the inverse discrete Fourier transform $\bm{p}^{(K)}$ of $\bm{\phi}^{(K)}$.
Return $\bm{p}^{(K)}$.
Algorithm 2 Computing the values of the pmf of $K$.
Proposition 3.3 of [Padmanabhan and Natarajan, 2021] provides another algorithm for computing the pmf of $K$, which although not relying on FFT, also involves an intertwined recursion. As mentioned in the Introduction, [Padmanabhan and Natarajan, 2021], referring to their model as a Bernoulli vector with a tree graphical model with conditional independence, employs the mean parameterization, and their approach to computing the pmf of $K$ is made possible by that.
4.2 Expected allocations
Other quantities the expression of the joint pgf allows to compute are those of the form $\mathbb{E}[J_{v}\mathbbm{1}_{\{K=k\}}]$, $v\in\mathcal{V}$, $k\in\mathbb{N}$, referred to as expected allocations. As highlighted in [Blier-Wong et al., 2025], such are key to many computations in risk management, for instance that of the conditional-mean risk-sharing rule of [Denuit and Dhaene, 2012] or of the contributions to the Tail-Value-at-Risk under Euler’s rule [Tasche, 2007]. For a further discussion in the case of the Ising model, see [Denuit and Robert, 2022]. [Blier-Wong et al., 2025] introduce the ordinary generating function of expected allocations (OGFEA) for efficient computations relying of the FFT algorithm.
Definition 4 (OGFEA).
Let $\bm{X}=(X_{v},\,v\in\mathcal{V})$ be a vector of random variables taking values in $\mathbb{N}^{d}$ and $Y$ be the sum of its components. The ordinary generating function of expected allocations (OGFEA) of $X_{v}$ to $Y$, $v\in\mathcal{V}$, is $\mathcal{P}_{Y}^{[v]}$ such that
$\mathcal{P}_{Y}^{[v]}(t)=\sum_{k=0}^{\infty}t^{k}\mathrm{E}[X_{v}\mathbbm{1}_{\{Y=k\}}],\quad t\in[-1,1],\,v\in\mathcal{V}.$
We derive the OGFEA of the Ising model from the joint pgf in Theorem 2. Note that here, $\mathrm{E}[J_{v}\mathbbm{1}_{\{K=k\}}]=\mathrm{Pr}(J=1,K=k)$.
Proposition 2.
Let $\bm{J}=(J_{v},\,v\in\mathcal{V})$ be a tree-structured Ising model defined on tree $T=(\mathcal{V},\mathcal{E})$, mean-parameterized with vectors of parameters $\bm{q}=(q_{v},\,v\in\mathcal{V})\in(0,1)^{d}$ and $\bm{\alpha}=(\alpha_{e},\,e\in\mathcal{E})\in\Lambda(\bm{q})$. Assume the underlying tree to be rooted in $j$, $j\in\mathcal{V}$. The OGFEA of $J_{j}$ to $K$ is given by
$\mathcal{P}_{K}^{[j]}(t)=q_{j}t\prod_{i\in\mathrm{ch}(j)}\xi_{i}^{\mathcal{T}_{j}}(t_{i\mathrm{dsc}(i)}),\quad t\in[-1,1],$
where $\xi_{v}^{\mathcal{T}_{j}}$, for $v\in\mathcal{V}$, is defined as in (16).
Proof.
As discussed above, the choice of the root bears no stochastic significance, and one may select any vertex for this purpose when deriving the joint pgf as in Theorem 2 without it affecting the result. Choosing $j$ as the root, we notice that $t_{j}$ is not in $\bm{t}_{i\mathrm{dsc}(i)}$ for all $i\in\mathrm{ch}(j)$. Therefore, the quantities $\prod_{i\in\mathrm{ch}(j)}\zeta_{i}^{\mathcal{T}_{j}}(\bm{t}_{i\mathrm{dsc}(i)})$ and $\prod_{i\in\mathrm{ch}(j)}\xi_{i}^{\mathcal{T}_{j}}(\bm{t}_{i\mathrm{dsc}(i)})$ are both simple coefficients to $t_{j}$. The result then follows from Theorem 2.4 of [Blier-Wong et al., 2025], stating that
[…]
As a pendant to Algorithm 2, the following algorithm indicates how to leverage the FFT algorithm to obtain the expected allocations for the Ising model.
Input: Vector of marginal parameters $\bm{q}=(q_{v})_{v\in\mathcal{V}}$, topologically ordered weighted adjacency matrix $\bm{A}=(A_{ij})_{i\times j\in\mathcal{V}\times\mathcal{V}}$.
[…]
Modify $\bm{A}$ to be topologically ordered according to vertex $v$. This can be done, for instance, by employing Algorithm 5 of [Côté et al., 2025].
Set $n_{\mathrm{FFT}}$ to be a large power of 2 (e.g. $2^{13}$). This determines $k_{\max}$.
Set $\bm{b}=(b_{i})_{i\in\{0,\ldots,n_{\mathrm{FFT}}-1\}}=(0,1,0,0,\ldots,0)$.
Use FFT to compute the discrete Fourier transform $\bm{\phi}^{(b)}$ of $\bm{b}$.
for $\ell=0,1,\dots,n_{\rm FFT}-1$ do
   Compute $\phi_{\ell}^{(K)}=\mathcal{P}_{J}(\phi_{\ell}^{(b)},\ldots,\phi_{\ell}^{(b)})$ by employing Algorithm 1, but change the penultimate step by “Compute $\mathcal{P}_{\bm{J}}=\xi^{(1)}$.”
Use fft to compute the inverse discrete Fourier transform $\bm{e}^{(K)}$ of $\bm{\phi}^{(K)}$.
Return $\bm{e}^{(K)}$.
[…]
5 Stochastic representation and sampling
In its most common formulations (3) and (8), the Ising model does not admit a straightforward sampling procedure as one cannot isolate the effect of one vertex on its neighbor. An implementation of Markov Chain Monte-Carlo, called the Gibbs sampler, circumvent this need. It produces a realization of an Ising model by iteratively simulating local interactions between random variables, refining the realization’s agreement to the joint distribution at each iteration. This, of course, requires many random numbers to produce a single realization of $\bm{J}$, and may not be efficient when $d$ is large. The Gibbs sampler method is described in details in Chapter 12 of [Koller and Friedman, 2009]. See [Izenman, 2021] for a complete exposition of the sampling problem for an Ising model. Most implementations of sampling methods for Ising models employ the Gibbs sampler or variations of it; for instance, the IsingSampler package in R [Epskamp, 2023]. An alternative to the Gibbs sampler is importance sampling, which consists of weighting realizations of a distribution close to the one from which we aim to sample and for whose a direct sampling method exists. It is more efficient than the Gibbs sampler, but comes with the drawback of accrued variance for Monte-Carlo estimations, see Chapter 12 of [Koller and Friedman, 2009].
Obviously, a direct sampling method, where one iteratively produces a realization of each component of $\bm{J}$ using a single random number, would be most efficient. This method is made possible for tree-structured Ising models by a a clique-based stochastic representation, provided in the next theorem. It is enabled by the use of the mean parameterization. Beyond providing a straightforward sampling procedure, the stochastic representation given in Theorem 3 is a novel result which deepens our understanding of the stochastic dynamics at play in a tree-structured Ising model. The stochastic representation illustrates clearly how the change in paradigm discussed in Section 2.4 affects one’s understanding of the model. For a further discussion on how the paradigm change in defining a MRF is reflected in the stochastic representation and sampling procedures, see Section 6 of [Côté et al., 2025].
Theorem 3.
For a tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$ and a chosen root $r\in\mathcal{V}$, let $\mathcal{T}_{r}$ be its rooted version. For vectors of parameters $\bm{q}=(q_{v},\,v\in\mathcal{V})\in[0,1]^{d}$ and $\bm{\alpha}=(\alpha_{e},\,e\in\mathcal{E})\in\Lambda(\bm{q})$, define $\bm{J}=(J_{v},\,v\in\mathcal{V})$ such that $J_{r}$ follows a Bernoulli distribution of parameter $q_{r}$ and, for $v\in\mathcal{V}\backslash\{r\}$, we have the construction
$J_{v}=(1-J_{\mathrm{pa}(v)})I_{v,0}^{\left\{q_{v}-\alpha_{(\mathrm{pa}(v),v)}\frac{\sigma_{(\mathrm{pa}(v),v)}}{1-q_{\mathrm{pa}(v)}}\right\}}+J_{\mathrm{pa}(v)}I_{v,1}^{\left\{q_{v}+\alpha_{(\mathrm{pa}(v),v)}\frac{\sigma_{(\mathrm{pa}(v),v)}}{q_{\mathrm{pa}(v)}}\right\}},$ (20)
[…]
Proof.
It is standard to check that $J_{v}$ is Bernoulli distributed with mean $q_{v}$ and that $\mathrm{Corr}(J_{\mathrm{pa}(v)},J_{v})=\alpha_{(\mathrm{pa}(v),v)}$, for every $v\in\mathcal{V}$. We argue that $\bm{J}$ is a MRF: For any $v\in\mathcal{V}$, the stochastic representation of $J_{v}$ depends only on three random variables $J_{\mathrm{pa}(v)}$, $I_{v,0}^{\{\beta_{v,0}\}}$ and $I_{v,1}^{\{\beta_{v,1}\}}$, the latter two being independent of the other idiosyncratic random variables of all other vertex $(I_{w,0}^{\{\beta_{w,0}\}},\,w\in\mathcal{V}\backslash\{v\})$ and $(I_{w,1}^{\{\beta_{w,1}\}},\,w\in\mathcal{V}\backslash\{v\})$. Therefore, dependence relations between random variables is solely induced through the filiation relation $\mathrm{pa}$. For $v\in\mathcal{V}$, dependence is encompassed by its parent random variable and those to whom it is a parent (its children) – that is, its neighbors. Knowing the value of all of its neighbors blocks the flow of dependence with random variables from any other vertex on $\mathcal{T}$ – and this means conditional independence with respect to the neighboring random variables. Thus, $\bm{J}$ satisfies the local Markov property (4) on $\mathcal{T}$ and is a MRF according to Definition 2. By Theorem 1, it is a tree-structured Ising model. ∎
The pmf associated to the stochastic representation in Theorem 3 is, of course, given by (12). As established before, the choice of the root allows for the filiation relations to be well-defined on the underlying tree, but has no incidence on the resulting distribution for $\bm{J}$.
As mentioned above, along with the clique-based stochastic representation in Theorem 3 comes a straightforward simulation algorithm. As for the previous algorithms, the adjacency matrix therein must be constructed in a topological order given a chosen root $r\in\mathcal{V}$. Filiation relations are deduced from such and provide the order of sampling for the components of $\bm{J}$. Again, that order depends on the chosen root $r$, but that choice has no distributional incidence.
Input: probability parameters $\bm{q}=(q_{v})_{v\in\mathcal{V}}$; topologically ordered weighted adjacency matrix $\bm{A}=(A_{ij})_{i\times j\in\mathcal{V}\times\mathcal{V}}$.
Output: Realization of $\bm{J}$.
[…]
for $v=2,\ldots,d$ do
   Compute $\pi_{v}=\inf\{j:A_{vj}>0\}$.
 if $J_{\pi_{v}}=0$ then
      Produce a realization of $J_{v}$ as a $\mathrm{Bernoulli}\left(q_{v}-A_{\pi_{v},v}\tfrac{\sigma_{(\mathrm{\pi_{v},v)}}}{1-q_{\pi_{v}}}\right)$ random variable.
 else
      Produce a realization of $J_{v}$ as a $\mathrm{Bernoulli}\left(q_{v}+A_{\pi_{v},v}\tfrac{\sigma_{(\pi_{v},v)}}{q_{\pi_{v}}}\right)$ random variable.
Return $(J_{v},\;v\in\{1,\ldots,d\})$.
Algorithm 4 Sampling from a tree-structured Ising model.
Under the fixed-conditional paradigm leading to the other, more common parameterizations, the conception of the stochastic dynamics is not clique-based but neighbor-based. Stochastic dynamics take into account every neighbor’s vertex, see [Besag, 1974]; they do not flow down the tree. A stochastic representation leaning on that could therefore not produce a straightforward sampling procedure as every vertex’s realization depends on the others’; hence the need for the Gibbs sampler.
[…]
$J_{v}=I^{\{\beta_{v}\}}_{v}J_{\mathrm{pa}(v)}+\left(1-I^{\{\beta_{v}\}}_{v}\right)(1-J_{\mathrm{pa}(v)}),\quad v\in\mathrm{dsc}(r),$ (21)
where $(I^{\{\beta_{v}\}}_{v},\,v\in\mathrm{dsc}(r))$ is a vector of independent Bernoulli trials of success probability $\beta_{v}$, and $\beta_{v}=0.5(\alpha_{(\mathrm{pa}(v),v)}+1)$ for all $v\in\mathrm{dsc}(r)$. To our knowledge, this is a novel stochastic representation of a tree-structured no-external-field Ising model. That the stochastic representation in (21) holds for this specific case may explain its simple mapping between the canonical parameterization and the mean parameterization; this ought to be explored further.
As discussed in [Fontana and Semeraro, 2018], all members of the Fréchet class of bivariate symmetric Bernoulli distributions lie on a segment within the convex hull of all bivariate Bernoulli distributions; for an illustration, see Figure 3 of [Blier-Wong et al., 2022]. Comonotonicity and counter-monotonicity [^4] Comonotonicity is perfect positive dependence. Counter-monotonicity is perfect negative dependence. are the extremities of that segment, and independence is its midpoint. This is why (21) captures all possible dependence schemes of tree-structured no-external-field Ising models, because it renders a discrete mixture between comonotonicity and counter-monotonicity for all of its pairwise distributions (as in (5)). In this extent, the stochastic representation described in Section 1.1 of [Nam et al., 2022], producing a discrete mixture between comonotonicity and independence, allows to describe all of those exhibiting positive dependence. This representation is understood as the broadcasting problem on trees in information theory. For edges $e$ with $\alpha_{e}<0$, one could conversely describe the stochastic dynamics as a mixture between independence and counter-monotonicity.
6 Poisson approximation
Poisson approximation refers to the well-known observation that a Poisson distribution becomes increasingly similar to a Bernoulli distribution as its mean parameter decreases closer to zero. It is not so much a computational tool for approximating either distribution by the other than a theoretical investigation of the distance between distributions as they tend to limit cases. In this section, we extend this observation to the multivariate case, for Ising models and a family of MRFs with Poisson marginal distributions.
Most theoretical works regarding Poisson approximations for dependent Bernoulli trials revolve around the Chen-Stein method from [Chen, 1975], notably discussed in [Arratia et al., 1990]; see [Barbour et al., 1992] and [Novak, 2019] for a review of such works. The Chen-Stein method assesses the error of approximating the sum of Bernoulli random variables, depending on their dependence schemes, by a singular Poisson random variable, which could be seen as the sum of independent Poisson random variables. The soundness of the approximation is contingent on the strength of the dependence within the Bernoulli random vector. Multivariate extensions of the Chen-Stein method (e.g. [Pianoforte and Turin, 2023], [Roos, 1999]) consider the sum of independent Bernoulli random vectors and approximates it by a singular Poisson random vector.
We proceed differently. We look at the similarity between vectors from two family of distribution, one with Bernoulli marginals (the Ising model) and the other with Poisson marginals (the family MPMRF, defined below). In other words, we add dependence between the Poisson random variables in order to mimic the dependence scheme of the Bernoulli trials in the Ising model. We proceed as such not only to have a multivariate outlook to the relation between Bernoulli and Poisson random variable; our intention is mostly to investigate whether the dependence schemes of these two fixed-marginal MRFs closely relate in terms of their strength and dynamics.
Denote by $\circ$ the binomial thinning operator introduced in [Steutel and van Harn, 1979]. For $\alpha\in(0,1)$ and a random variable $X$ taking values in $\mathbb{N}$, binomial thinning is the operation $\alpha\circ X=\sum_{i=1}^{X}I^{\{\alpha\}}_{i}$, where $\{I^{\{\alpha\}}_{i},\,i\in\mathbb{N}\}$ is a sequence of independent Bernoulli trials of success probability $\alpha$.
Definition 5 (MPMRF).
For a tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$, choose a root $r\in\mathcal{V}$ and let $\mathcal{T}_{r}$ be its rooted version. Given parameters $\lambda>0$ and $\bm{\alpha}=(\alpha_{e},\,e\in\mathcal{E})\in(0,1)^{d-1}$, let $\bm{N}=(N_{v},\,v\in\mathcal{V})$ be a random vector defined according to the following stochastic representation
$N_{v}=\begin{cases}L_{r},&\text{if }v=r;\\ \alpha_{(\mathrm{pa}(v),v)}\circ N_{\mathrm{pa}(v)}+L_{v},&\text{if }v\in dsc(r)\end{cases},\quad v\in\mathcal{V},$ (22)
[…]
By Theorem 2 of [Côté et al., 2025], the choice of the root has no incidence on the distribution of the vector $\bm{N}$, hence why we do not specify it for its distributional notation. We want to assess the closeness of the distributions of $\bm{N}$ and $\bm{J}$ when marginal distributions tend to degenerate at 0. This intent is greatly facilitated by the mean parameterization: we directly control the marginal distributions via the vector of parameter $\bm{q}$, thus our observations may be made by letting $\bm{q}$ be small, simply.
The total variation distance between two $d$-dimensional vectors of random variables, say $\bm{X}$ and $\bm{Y}$, is given by
$\mathbbm{d}_{\mathrm{TV}}(\bm{X},\bm{Y})=\sup_{A\in\mathcal{B}(\mathbb{R}^{d})}\left\{|\mathrm{Pr}(\bm{X}\in A)-\mathrm{Pr}(\bm{Y}\in A)|\right\},$
where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel set of $\mathbb{R}^{d}$. The total variation distance may be employed to measure the closeness of two multivariate distributions.
Theorem 4.
[…]
Proof.
The proof is provided in Appendix A.2. ∎
The approximation given by Theorem 4 becomes better as $q$ diminishes not just because the scale of both distributions decreases: the bound obtained on $\mathbbm{d}_{TV}(\bm{J},\bm{N})$ is of magnitude $O(q^{2})$ and not just $O(q)$. Note that Theorem 4 holds for any vector of dependence parameters $\bm{\alpha}$ and that the bound for $\mathbbm{d}_{\rm TV}(\bm{J},\bm{N})$ is not affected by its value. The dependence between the elements of $\bm{J}$ may be strong or not, the distribution of $\bm{N}$ is a good approximation if $q$ is small.
As a special case of Theorem 4, we have that a Markov-Bernoulli distribution may be closely approximated by a Poisson AR(1) distribution as the marginal parameter of each becomes small, given they are the respective distributions of $\bm{J}$ and $\bm{N}$ when defined on a series tree. Poisson AR(1) distributions were introduced in [McKenzie, 1985]; for a comment on their connection to MPMRF, see Section 2 of [Côté et al., 2025]. A direct consequence of Theorem 4 is that the distributions of the sum of each random vector also grow closer when $q$ becomes small, as made explicit in the following corollary.
Corollary 1.
Under the settings of Theorem 4, as $q$ diminishes, the distribution of $M=\sum_{v\in\mathcal{V}}N_{v}$ increasingly becomes a good approximation of the distribution of $K=\sum_{v\in\mathcal{V}}J_{v}$, for $\mathbbm{d}_{\mathrm{TV}}(K,M)\leq 1.2dq^{2}$.
Proof.
Since the set $\{\bm{x}\in\mathbb{R}^{d}:\sum_{v=1}^{d}x_{v}\in C\}$, for any $C\in\mathcal{B}(\mathbb{N})$, is an element of $\mathcal{B}(\mathbb{N}^{d})$, then
$\sup_{C\in\mathcal{B}(\mathbb{N})}\left\{|\mathrm{Pr}(K\in C)-\mathrm{Pr}(M\in C)|\right\}\leq\sup_{A\in\mathcal{B}(\mathbb{N}^{d})}\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in A)|\right\}.$
Hence, $\mathbbm{d}_{\mathrm{TV}}(K,M)\leq 1.2dq^{2}$ given Theorem 4. ∎
As we show in the next theorem, the variability (or riskiness) of $M$ always remains greater than that of $K$, making the Poisson approximation a conservative approximation for a risk modeler. This comparison is made with the convex order, defined as follows: A random variable $Y$ dominates another random variable $X$ with respect to the convex order, noted $X\preceq_{\rm cx}Y$, if $\mathbb{E}[\varphi(X)]\leq\mathbb{E}[\varphi(Y)]$ for all convex functions $\varphi$ such that both expectations are well-defined. In economics and actuarial science, orderings in terms of the convex order mean the embodiment of more risk in distributions, see [Rothschild and Stiglitz, 1978]. Textbooks [Müller and Stoyan, 2002] and [Shaked and Shanthikumar, 2007] offer an introduction to the convex order and to integral stochastic orders in general.
Theorem 5.
Under the settings of Theorem 4, $K\preceq_{cx}M$.
Proof.
The proof is provided in Appendix A.3. ∎
The Poisson approximation presented in the previous results deepens our understanding of Ising models by connecting the two families of fixed-marginal MRFs. For instance, as shown by the representation in (22), the distribution of $\bm{N}$ is predicated on a propagation-innovation stochastic dynamic. This comes with nice interpretations for applications as such dynamics are common for network-based phenomena. This behavior does not apply to the Ising model, however, as shown by the stochastic representation in Theorem 3, but the Poisson approximation shows that a propagation-based interpretation of an Ising model’s dynamics would still be valid for rare events on a tree.
Another aspect of the theoretical interest of the above results comes in connection with [Côté et al., 2024]. In this research article, the authors analyzes topological properties of the family MPMRF. They answer questions such as which component of $\bm{N}$ has more influence on $M$, depending on the position of its vertex within the tree, or which underlying tree’s shape yield a riskier distribution of $M$. The proof of the cornerstone result of [Côté et al., 2024], Theorem 2, enabling the considerations regarding vertex centrality and tree-shape orderings for the model, relies on the self-decomposability of the Poisson distribution and thus its approach cannot be transposed to the Ising model. The closeness of the joint distributions of the two models given by Theorem 4 signals that these results nonetheless extend to the Ising model, at least when $q$ is small enough.
7 Numerical example
In this example, we showcase the results of Sections 4, 5 and 6. We produce the exact values obtained using the joint pgf and FFT method, relying on Algorithm 2, and values obtained with Monte-Carlo methods, relying on Algorithm 4 to produce samples of $\bm{J}$. We want to show the relevance of the Poisson approximation: we compute values of the pmf of $K$ and examine how well the pmf of $M$ approximates its values. The values of the pmf of $M$ are obtained by using Algorithm 3 of [Côté et al., 2025].
Let $\bm{J}=(J_{1},\ldots,J_{7})$ be a vector of random variables following an Ising model with its underlying tree being a binary tree of radius 2. We suppose $\alpha_{e}=0.7$ for every $e\in\mathcal{E}$. Results employing the are rendered in Table 1. As expected given Theorem 1, the values provided by the Poisson approximation are within $1.2dq^{2}=0.00084$ of the exact value, and even suggest that the bound on $\mathbbm{d}_{\mathrm{TV}}(K,M)$ could probably be tighter. We also remark that the Poisson approximation falls within the 90% confidence intervals for the values obtained from Monte-Carlo methods when using $n\in\{1000,10000\}$ samples. These confidence intervals were derived by repeating Monte-Carlo estimation 1000 times.
$\begin{subarray}{c}\text{Exact}\\ \text{value}\end{subarray}$ $\begin{subarray}{c}\text{Poisson}\\ \text{approximation}\end{subarray}$ $\begin{subarray}{c}\text{Monte-Carlo IC90\%}\\ (n=1000)\end{subarray}$ $\begin{subarray}{c}\text{Monte-Carlo IC90\%}\\ (n=10000)\end{subarray}$
$\mathrm{Pr}(K=0)$ 0.97231 0.97239 [0.963, 0.981] [0.9697, 0.9751]
$\mathrm{Pr}(K=1)$ 0.01309 0.01307 [0.007, 0.019] [0.0111, 0.0149]
$\mathrm{Pr}(K=2)$ 0.00289 0.00291 [0.000, 0.006] [0.0021, 0.0037]
$\mathrm{Pr}(K\geq 3)$ 0.01170 0.01164 [0.006, 0.017] [0.0100, 0.0135]
Table 1: Comparison of the values of the pmf of $K$, the sum of the components of an Ising model with $q=0.01$, using different computation methods
Consider the function $\pi_{X}:\mathbb{R}\mapsto\mathbb{R}$ given by $\pi_{X}(z)=\mathrm{E}[(X-z)_{+}]$, for $z\in\mathbb{R}$. For random variables $X$ and $Y$, we have $X\preceq_{\rm cx}Y$ if and only if $\pi_{X}(z)\leq\pi_{Y}(z)$ for all $z\in\mathbb{R}$ and $\mathrm{E}[X]=\mathrm{E}[Y]$, see Proposition 3.4.3 of [Denuit et al., 2006]. One could be interested in the values taken by $\pi_{K}$ and $\pi_{M}$ to confirm the convex order between $K$ and $M$ as stated in Theorem 5. The values should be very similar, yet those of $M$ should always be superior. Table 2 confirms this.
$\begin{subarray}{c}\text{Exact}\\ \text{value}\end{subarray}$ $\begin{subarray}{c}\text{Poisson}\\ \text{approximation}\end{subarray}$
$\pi_{K}(0)$ 0.07000 0.07000
$\pi_{K}(1)$ 0.04231 0.04239
$\pi_{K}(2)$ 0.02772 0.02785
$\pi_{K}(3)$ 0.01602 0.01621
$\pi_{K}(4)$ 0.00901 0.00922
$\pi_{K}(5)$ 0.00446 0.00466
$\pi_{K}(6)$ 0.00121 0.00141
$\pi_{K}(7)$ 0.00000 0.00016
Table 2: Values of functions of $\pi_{K}$, an Ising model with $q=0.01$, and that of its Poisson approximation.
[…]
$\begin{subarray}{c}\text{Exact}\\ \text{value}\end{subarray}$ $\begin{subarray}{c}\text{Poisson}\\ \text{approximation}\end{subarray}$ $\begin{subarray}{c}\text{Monte-Carlo IC90\%}\\ (n=1000)\end{subarray}$ $\begin{subarray}{c}\text{Monte-Carlo IC90\%}\\ (n=10000)\end{subarray}$
$\mathrm{Pr}(K=0)$ 0.9972031 0.9972039 [0.994, 1.000] [0.9963, 0.9981]
$\mathrm{Pr}(K=1)$ 0.0013405 0.0013402 [0.000, 0.004] [0.0008, 0.0019]
$\mathrm{Pr}(K=2)$ 0.0002897 0.0002899 [0.000, 0.001] [0.0001, 0.0006]
$\mathrm{Pr}(K\geq 3)$ 0.0011666 0.0011660 [0.000, 0.003] [0.0006, 0.0017]
Table 3: Comparison of the values of the pmf of $K$, the sum of the components of an Ising model with $q=0.001$, using different computation methods
The similitude between functions $\pi_{K}$ and $\pi_{M}$ is also increased, with that of the Poisson approximation being greater only by a jot. This is seen in Table 4.
$\begin{subarray}{c}\text{Exact}\\ \text{value}\end{subarray}$ $\begin{subarray}{c}\text{Poisson}\\ \text{approximation}\end{subarray}$
$\pi_{K}(0)$ 0.007000 0.007000
$\pi_{K}(1)$ 0.004203 0.004204
$\pi_{K}(2)$ 0.002747 0.002748
$\pi_{K}(3)$ 0.001580 0.001582
$\pi_{K}(4)$ 0.000888 0.000890
$\pi_{K}(5)$ 0.000438 0.000440
$\pi_{K}(6)$ 0.000118 0.000120
$\pi_{K}(7)$ 0.000000 0.000002
Table 4: Values of functions of $\pi_{K}$, an Ising model with $q=0.01$, and that of its Poisson approximation.
Acknowledgment
This work was partially supported by the Natural Sciences and Engineering Research Council of Canada (Cossette: 04273; Marceau: 05605; Côté: 581589859) and by the Chaire en actuariat de l’Université Laval (Marceau).
Appendix A Proofs
A.1 Proof of Proposition 1
The case where $(u,v)\in\mathcal{E}$ is trivial due to the mean parameterization. We prove for the general case where $k\geq 3$ vertices are on $\mathrm{path}(u,v)$. Let us denote the vertices on the path by $\{\ell_{i},\;i\in\{1,\ldots,k\}\}$ such that $\ell_{1}=u$ and $\ell_{k}=v$. Hence, $\mathrm{Cov}(J_{u},J_{v})=\mathrm{Cov}(J_{\ell_{1}},J_{\ell_{v}})$. By conditioning on $J_{\ell_{k-1}}$,
$\mathrm{Cov}(J_{\ell_{1}},J_{\ell_{k}})=\mathrm{Cov}(\mathrm{E}[J_{\ell_{1}}|J_{\ell_{k\text{-}1}}],\mathrm{E}[J_{\ell_{k}}|J_{\ell_{k\text{-}1}}])+\mathrm{E}[\mathrm{Cov}(J_{\ell_{1}},J_{\ell_{k}}|J_{\ell_{k\text{-}1}})].$ (23)
The second term in (23) is equal to zero by the global Markov property. Note that, from the mean parameterization fixing bivariate distributions associated to edges, for any $m\in\{2,\ldots,k\}$,
$\mathrm{E}[J_{\ell_{m}}|J_{\ell_{m\text{-}1}}]$ $=\left(1-J_{\ell_{m\text{-}1}}\right)\left(q_{\ell_{m}}-\frac{\alpha_{(\ell_{m\text{-}1},\ell_{m})}\sigma_{(\ell_{m\text{-}1},\ell_{m})}}{1-q_{\ell_{m\text{-}1}}}\right)+J_{\ell_{m\text{-}1}}\left(q_{\ell_{m}}+\frac{\alpha_{(\ell_{m\text{-}1},\ell_{m})}\sigma_{(\ell_{m\text{-}1},\ell_{m})}}{q_{\ell_{m\text{-}1}}}\right)$
$=q-\frac{\alpha_{(\ell_{m\text{-}1},\ell_{m})}\sigma_{(\ell_{m\text{-}1},\ell_{m})}}{1-q_{\ell_{m\text{-}1}}}+\frac{\alpha_{(\ell_{m\text{-}1},\ell_{m})}\sigma_{(\ell_{m\text{-}1},\ell_{m})}}{q_{\ell_{m\text{-}1}}(1-q_{\ell_{m\text{-}1})}}J_{\ell_{m\text{-}1}}.$
Therefore, (23) becomes
$\mathrm{Cov}(J_{\ell_{1}},J_{\ell_{k}})$ $=\mathrm{Cov}\left(\mathrm{E}[J_{\ell_{1}}|J_{\ell_{k\text{-}1}}],q-\frac{\alpha_{(\ell_{k\text{-}1},\ell_{k})}\sigma_{(\ell_{k\text{-}1},\ell_{k})}}{1-q_{\ell_{k\text{-}1}}}+\frac{\alpha_{(\ell_{k\text{-}1},\ell_{k})}\sigma_{(\ell_{k\text{-}1},\ell_{k})}}{q_{\ell_{k\text{-}1}}(1-q_{\ell_{k\text{-}1})}}J_{\ell_{k\text{-}1}}\right)$
$=\frac{\alpha_{(\ell_{k\text{-}1},\ell_{k})}\sigma_{(\ell_{k\text{-}1},\ell_{k})}}{q_{\ell_{k\text{-}1}}(1-q_{\ell_{k\text{-}1})}}\mathrm{Cov}\left(\mathrm{E}[J_{\ell_{1}}|J_{\ell_{k\text{-}1}}],J_{\ell_{k\text{-}1}}\right).$
Repeating this rationale iteratively, we obtain
$\mathrm{Cov}(J_{\ell_{1}},J_{\ell_{k}})=\left(\prod_{m=2}^{k}\frac{\alpha_{(\ell_{k\text{-}1},\ell_{k})}\sigma_{(\ell_{k\text{-}1},\ell_{k})}}{q_{\ell_{k\text{-}1}}(1-q_{\ell_{k\text{-}1})}}\right)\mathrm{Var}(J_{\ell_{1}})$ $=\left(\prod_{m=2}^{k}\alpha_{(\ell_{k\text{-}1},\ell_{k})}\right)\left(\frac{\sigma_{(\prod_{m=2}^{k}\ell_{k\text{-}1},\ell_{k})}}{\prod_{j=2}^{k-1}q_{\ell_{j}(1-q_{\ell_{j\text{+}1})}}}\right)$
$=\left(\prod_{m=2}^{k}\alpha_{(\ell_{k\text{-}1},\ell_{k})}\right)\sqrt{q_{l_{1}}(1-q_{l_{1}})q_{l_{k}}(1-q_{l_{k}})}.$ (24)
Reverting back to the original indices, (24) is rewritten as
$\mathrm{Cov}(J_{u},J_{v})=\left(\prod_{e\in\mathrm{path}(u,v)}\alpha_{e}\right)\sqrt{q_{u}(1-q_{u})q_{v}(1-q_{v})}.$
The desired result directly ensues. $\square$
A.2 Proof of Theorem 4
Let $\mathcal{B}^{\ast}(\Omega^{d}):=((A_{i},i\in\{1,\ldots,d\})\in\mathcal{B}(\Omega^{d}):A_{i}\neq\varnothing,i\in\{1,\ldots,d\})$, for a one-dimensional space set $\Omega$, i.e. $\mathcal{B}^{\ast}(\Omega^{d})$ is the Borel set $\mathcal{B}(\Omega^{d})$ minus its elements for which one or more components are the empty set. Because $\mathrm{Pr}(\bm{J}\in A)>0$ only if $A\in\mathcal{B}^{\ast}(\{0,1\}^{d})$ and $\mathrm{Pr}(\bm{N}\in A)>0$ only if $A\in\mathcal{B}^{\ast}(\mathbb{N}^{d})$, we have
$\mathbbm{d}_{\mathrm{TV}}(\bm{J},\bm{N})$ $=\sup_{A\in\mathcal{B}^{\ast}(\mathbb{N}^{d})}\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in A)|\right\}$
$=\max\left(\sup_{A\in\mathcal{B}^{\ast}(\{0,1\}^{d})}\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in A)|\right\};\sup_{A\in\mathcal{B}^{\ast}(\mathbb{N}^{d})\backslash\mathcal{B}^{\ast}(\{0,1\}^{d})}\left\{\mathrm{Pr}(\bm{N}\in A)\right\}\right)$ (25)
Let $\{B_{i},\;i\in\{0,\ldots,d\}\}$ be a partition of ($\mathcal{B}^{\ast}(\{0,1\}^{d})$, defined such that $B_{i}$ comprises the elements of ($\mathcal{B}^{\ast}(\{0,1\}^{d})$ with $i$ components being a singleton. We study the first supremum in (25). For that matter, we define the isomorphism $\psi:\{0,1\}^{d}\to\mathbb{N}^{d}$ such that
$\psi(A)_{v}=\left\{\begin{array}[]{ll}A_{v}&\text{if }A_{v}\subset\{0,1\}\\ \mathbb{N}&\text{if }A_{v}=\{0,1\}\end{array}\right.,\quad A\in\mathcal{B}^{\ast}(\{0,1\}).$
We then note
$\sup_{A\in\mathcal{B}^{\ast}(\{0,1\}^{d})}{\scriptstyle\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in A)|\right\}}$
$\quad=\max_{i\in\{0,\ldots,d\}}\left(\sup_{A\in B_{i}}{\scriptstyle\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in A)|\right\}}\right)$
$\quad\leq\max_{i\in\{0,\ldots,d\}}\left(\sup_{A\in B_{i}}{\scriptstyle\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in\psi(A))|\;+\;|\mathrm{Pr}(\bm{N}\in\psi(A))-\mathrm{Pr}(\bm{N}\in A)|\right\}}\right)$
$\quad\leq\max_{i\in\{0,\ldots,d\}}\left(\sup_{A\in B_{i}}{\scriptstyle\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in\psi(A))|\right\}}+\sup_{A\in B_{i}}{\scriptstyle\left\{|\mathrm{Pr}(\bm{N}\in\psi(A))-\mathrm{Pr}(\bm{N}\in A)|\right\}}\right).$ (26)
On one hand, by the conditional independence properties of both models, we note, for $i\in\{0,\ldots,d\}$,
$\sup_{A\in B_{i}}{\scriptstyle\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in\psi(A))|\right\}}$
$=\sup_{A\in B_{i}}{\scriptstyle\left\{|\mathrm{Pr}(J_{r}\in A_{r})\prod_{v\in\mathcal{V}\backslash\{r\}}\mathrm{Pr}(J_{v}\in A_{v}|J_{\mathrm{pa}(v)}\in A_{\mathrm{pa}(v)})-\mathrm{Pr}(N_{r}\in\psi(A)_{r})\prod_{v\in\mathcal{V}\backslash\{r\}}\mathrm{Pr}(N_{v}\in\psi(A)_{v}|N_{\mathrm{pa}(v)}\in\psi(A)_{\mathrm{pa}(v)})|\right\}}$
$=\sup_{A\in B_{i}}\left\{\left|\mathrm{Pr}(J_{r}\in A_{r})\left(\prod_{\begin{subarray}{c}v\in\mathcal{V}\\ A_{v}\in\{0,1\}\\ A_{\mathrm{pa}(v)}\in\{0,1\}\end{subarray}}\mathrm{Pr}(J_{v}=A_{v}|J_{\mathrm{pa}(v)}=A_{\mathrm{pa}(v)})\right)\left(\prod_{\begin{subarray}{c}v\in\mathcal{V}\\ A_{v}\in\{0,1\}\\ A_{\mathrm{pa}(v)}=\{0,1\}\end{subarray}}\mathrm{Pr}(J_{v}=A_{v})\right)\right.\right.$
$\quad-\left.\left.\mathrm{Pr}(N_{r}\in\psi(A)_{r})\left(\prod_{\begin{subarray}{c}v\in\mathcal{V}\\ A_{v}\in\{0,1\}\\ A_{\mathrm{pa}(v)}\in\{0,1\}\end{subarray}}\mathrm{Pr}(N_{v}=A_{v}|N_{\mathrm{pa}(v)}=A_{\mathrm{pa}(v)})\right)\left(\prod_{\begin{subarray}{c}v\in\mathcal{V}\\ A_{v}\in\{0,1\}\\ A_{\mathrm{pa}(v)}=\{0,1\}\end{subarray}}\mathrm{Pr}(N_{v}=A_{v})\right)\right|\right\}.$ (27)
Indeed, for any $A\in B_{i}$, $i\in\{0,\ldots,1\}$, $A_{v}$ is either a singleton or the support of $J_{v}$; similarly, $\psi(A)_{v}$ is either a singleton or the support of $N_{v}$, $v\in\mathcal{V}$.
Firstly, let $\varphi_{v}(a):=|\scriptstyle{\mathrm{Pr}(J_{v}=a)-\mathrm{Pr}(N_{v}=a)}|$. We observe, for any $v\in\mathcal{V}$,
$\varphi_{v}(0)$ $=|1-q-\mathrm{e}^{-q}|\leq q^{2}$
$\varphi_{v}(1)$ $=|q-q\mathrm{e}^{-q}|\leq q^{2}.$
Therefore, $\varphi_{v}(a)\leq q^{2}$, $a\in\{0,1\}$.
Secondly, let $\widetilde{\varphi}_{v}(a):=|\scriptstyle{\mathrm{Pr}(J_{v}=a|J_{\mathrm{pa}(v)}=0)-\mathrm{Pr}(N_{v}=a|N_{\mathrm{pa}(v)}=0)}|$. We observe, for any $v\in\mathcal{V}\backslash\{r\}$,
$\widetilde{\varphi}_{v}(0)$ $=\left|(1-q(1-\alpha_{(\mathrm{pa}(v),v)}))-\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}\right|\leq q^{2}$
$\widetilde{\varphi}_{v}(1)$ $=\left|(q-\alpha_{(\mathrm{pa}(v),v)}q)-q(1-\alpha_{(\mathrm{pa}(v),v)})\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}\right|\leq q^{2}$
Hence, $\widetilde{\varphi}_{v}(a)\leq q^{2}$, $a\in\{0,1\}$.
Thirdly, let $\breve{\varphi}_{v}(a):=|\scriptstyle{\mathrm{Pr}(J_{\mathrm{pa}(v)}=1)\mathrm{Pr}(J_{v}=a|J_{\mathrm{pa}(v)}=1)-\mathrm{Pr}(N_{\mathrm{pa}(v)}=1)\mathrm{Pr}(N_{v}=a|N_{\mathrm{pa}(v)}=1)}|$. We observe, for any $v\in\mathcal{V}\backslash\{r\}$,
$\breve{\varphi}_{v}(0)$ $=\left|q(1-\alpha_{(\mathrm{pa}(v),v)}-q+q\alpha_{(\mathrm{pa}(v),v)})-q\mathrm{e}^{-q}(1-\alpha_{(\mathrm{pa}(v),v)})\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}\right|\leq q^{2}$
$\breve{\varphi}_{v}(1)$ $=\left|q(\alpha_{(\mathrm{pa}(v),v)}+q-\alpha_{(\mathrm{pa}(v),v)}q)-q\mathrm{e}^{-q}\left(\alpha_{(\mathrm{pa}(v),v)}\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}+q(1-\alpha_{(\mathrm{pa}(v),v)})^{2}\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}\right)\right|\leq 1.2q^{2}$
Thus, $\breve{\varphi}_{v}(a)\leq 1.2q^{2}$, $a\in\{0,1\}$.
Fourthly, let $\ddot{\varphi}_{v}(a):=|\scriptstyle{\mathrm{Pr}(J_{\mathrm{pa}(v)}=1|J_{\mathrm{pa}(\mathrm{pa}(v))}=0)\mathrm{Pr}(J_{v}=a|J_{\mathrm{pa}(v)}=1)-\mathrm{Pr}(N_{\mathrm{pa}(v)}=1|N_{\mathrm{pa}(\mathrm{pa}(v))}=0)\mathrm{Pr}(N_{v}=a|N_{\mathrm{pa}(v)}=1)}|$. We observe, for any $v\in\mathcal{V}\backslash(\{r\}\cup\mathrm{ch}(r))$,
$\ddot{\varphi}_{v}(0)$ $=\left|\scriptstyle{(q-\alpha_{(\mathrm{pa}(\mathrm{pa}(v)),\mathrm{pa}(v))}q)(1-\alpha_{(\mathrm{pa}(v),v)}-q+q\alpha_{(\mathrm{pa}(v),v)})}\right.$
$\quad\quad-\left.\scriptstyle{q(1-\alpha_{(\mathrm{pa}(\mathrm{pa}(v)),\mathrm{pa}(v))})\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(\mathrm{pa}(v)),\mathrm{pa}(v))})}(1-\alpha_{(\mathrm{pa}(v),v)})\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}}\right|\leq q^{2}$
$\ddot{\varphi}_{v}(1)$ $=\left|\scriptstyle{(q-\alpha_{(\mathrm{pa}(\mathrm{pa}(v)),\mathrm{pa}(v))}q)(\alpha_{(\mathrm{pa}(v),v)}+q-\alpha_{(\mathrm{pa}(v),v)}q)}\right.$
$\quad\quad-\left.\scriptstyle{q(1-\alpha_{(\mathrm{pa}(\mathrm{pa}(v)),\mathrm{pa}(v))})\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(\mathrm{pa}(v)),\mathrm{pa}(v))})}\left(\alpha_{(\mathrm{pa}(v),v)}\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}+q(1-\alpha_{(\mathrm{pa}(v),v)})^{2}\mathrm{e}^{-q(1-\alpha_{(\mathrm{pa}(v),v)})}\right)}\right|\leq 1.2q^{2}$
Consequently, $\ddot{\varphi}_{v}(a)\leq 1.2q^{2}$, $a\in\{0,1\}$.
We have the following inequality, for $a_{1},a_{2},b_{1},b_{2},b_{3},b_{4}\in[0,1]$,
$|a_{1}b_{1}b_{2}-a_{2}b_{3}b_{4}|\leq|a_{1}(b_{1}+b_{2})-a_{2}(b_{3}+b_{4})|\leq|a_{1}b_{2}-a_{2}b_{3}|+|a_{1}b_{2}-a_{2}b_{4}|.$
Therefore, each element considered for the supremum in (27) is dominated by a sum of at most $i$ among values of $\varphi$, $\widetilde{\varphi}$, $\breve{\varphi}$ or $\ddot{\varphi}$, all of which are dominated by $1.2q^{2}$. For example, considering the models encrypted on a three-vertex series tree and $r=1$, for $A=(\{1\},\{1\},\{1\})$,
$|\scriptstyle{\mathrm{Pr}(J_{1}=1)\mathrm{Pr}(J_{2}=1|J_{1}=1)\mathrm{Pr}(J_{3}=1|J_{2}=1)-\mathrm{Pr}(N_{1}=1)\mathrm{Pr}(N_{2}=1|N_{1}=1)\mathrm{Pr}(N_{3}=1|N_{2}=1)}|$
$\quad\leq|\scriptstyle{\mathrm{Pr}(J_{1}=1)\mathrm{Pr}(J_{2}=1|J_{1}=1)-\mathrm{Pr}(N_{1}=1)\mathrm{Pr}(N_{2}=1|N_{1}=1)|+|\mathrm{Pr}(J_{1}=1)\mathrm{Pr}(J_{3}=1|J_{2}=1)-\mathrm{Pr}(N_{1}=1)\mathrm{Pr}(N_{3}=1|N_{2}=1)}|$
$\quad=\breve{\varphi}_{2}(1)+\breve{\varphi}_{3}(1)$
since $J_{1}$ and $J_{2}$ have the same marginal distributions and likewise for $N_{1}$ and $N_{2}$. Each element of the supremum in (27) is thus dominated by $1.2iq^{2}$: that supremum is therefore dominated by $1.2iq^{2}$.
On the other hand, we observe, for any $A\in B_{i}$, $i\in\{0,\ldots,d\}$,
$\mathrm{Pr}(\bm{N}\in\psi(A))-\mathrm{Pr}(\bm{N}\in A)=\mathrm{Pr}\left(\bigcup_{v\in\mathcal{V}}N_{v}\in\psi(A)_{v}\backslash A_{v}\right)\leq\sum_{\begin{subarray}{c}v\in\mathcal{V}\\ A_{v}=\{0,1\}\end{subarray}}\mathrm{Pr}(N_{v}\geq 2).$
[…]
$\sum_{v\in\mathcal{V}}\mathrm{Pr}(N_{v}\geq 2)\leq(d-i)\left(1-e^{-q}-qe^{-q}\right)\leq(d-i)q^{2}$
Hence, $\sup_{A\in\mathcal{B}^{\ast}(\{0,1\}^{d})}{\scriptstyle\left\{|\mathrm{Pr}(\bm{N}\in\psi(A))-\mathrm{Pr}(\bm{N}\in A)|\right\}}\leq(d-i)q^{2}$. This, combined with (27)’s being dominated by $1.2iq^{2}$, yields, given (26),
$\sup_{A\in\mathcal{B}^{\ast}(\{0,1\}^{d})}{\scriptstyle\left\{|\mathrm{Pr}(\bm{J}\in A)-\mathrm{Pr}(\bm{N}\in A)|\right\}}\leq\max_{i\in\{0,\ldots,d\}}\{1.2iq^{2}+(d-i)q^{2}\}\leq 1.2dq^{2}.$ (28)
Next, we study the second supremum in (25); we note
$\sup_{A\in\mathcal{B}^{\ast}(\mathbb{N}^{d})\backslash\mathcal{B}^{\ast}(\{0,1\}^{d})}\left\{\mathrm{Pr}(\bm{N}\in A)\right\}$ $=\mathrm{Pr}(\max(\bm{N})\geq 2)$
$\leq\sum_{v\in\mathcal{V}}\mathrm{Pr}(N_{v}\geq 2)$
$=d\left(1-e^{-q}-qe^{-q}\right)\leq dq^{2}.$ (29)
Combining (25), (28) and (29) leads to
$\mathbbm{d}_{\mathrm{TV}}(\bm{J},\bm{N})\leq\max(1.2dq^{2},dq^{2})=1.2dq^{2}$
That is, as $q$ approaches 0, the total variation distance between $\bm{J}$ and $\bm{N}$ abates at the same rate as $q^{2}$. $\square$
A.3 Proof of Theorem 5
Let $N_{v}^{\prime}:=\alpha_{(\mathrm{pa}(v),v)}\circ J_{\mathrm{pa}(v)}+L_{v}$, $v\in\mathcal{V}\backslash\{r\}$. Thus $N_{v}^{\prime}$ defines itself on a hybrid of both models; its purpose is to allow working iteratively on vertices during the proof while the model remains coherent and well-defined. First, we prove $(J_{v}|J_{\mathrm{pa}(v)=x_{\mathrm{pa}(v)}})\preceq_{cx}(N_{v}^{\prime}|J_{\mathrm{pa}(v)=x_{\mathrm{pa}(v)}})$, $v\in\mathcal{V}\backslash\{r\}$, $x_{\mathrm{pa}(v)}\in\{0,1\}$. For a start, we have
$\mathrm{Pr}(J_{v}=0|J_{\mathrm{pa}(v)}=0)=1-(1-\alpha_{(\mathrm{pa}(v),v)})q\leq\mathrm{e}^{-(1-\alpha_{(\mathrm{pa}(v),v)})q}=\mathrm{Pr}(N_{v}^{\prime}=0|J_{\mathrm{pa}(v)}=0)$
$\mathrm{Pr}(J_{v}\leq 1|J_{\mathrm{pa}(v)}=0)=1\geq(1+(1-\alpha_{(\mathrm{pa}(v),v)})q)\mathrm{e}^{-(1-\alpha_{(\mathrm{pa}(v),v)})q}=\mathrm{Pr}(N_{v}^{\prime}\leq 1|J_{\mathrm{pa}(v)}=0),$
which implies $(J_{v}|J_{\mathrm{pa}(v)}=0)\preceq_{cx}(N_{v}^{\prime}|J_{\mathrm{pa}(v)}=0)$ by the Karlin-Novikoff sufficient condition for convex ordering and the fact that both sides of the order relation have the same mean. Similarly,
$\mathrm{Pr}(J_{v}=0|J_{\mathrm{pa}(v)}=1)=(1-\alpha_{(\mathrm{pa}(v),v)})(1-q)\leq(1-\alpha_{(\mathrm{pa}(v),v)})\mathrm{e}^{-(1-\alpha_{(\mathrm{pa}(v),v)})q}=\mathrm{Pr}(N_{v}^{\prime}=0|J_{\mathrm{pa}(v)}=1)$
$\mathrm{Pr}(J_{v}\leq 1|J_{\mathrm{pa}(v)}=1)=1\geq(\alpha_{(\mathrm{pa}(v),v)}+(1-\alpha_{(\mathrm{pa}(v),v)})^{2}q)\mathrm{e}^{-(1-\alpha_{(\mathrm{pa}(v),v)})q}=\mathrm{Pr}(N_{v}^{\prime}\leq 1|J_{\mathrm{pa}(v)}=1),$
which implies $(J_{v}|J_{\mathrm{pa}(v)}=1)\preceq_{cx}(N_{v}^{\prime}|J_{\mathrm{pa}(v)}=1)$ again by Karlin-Novikoff and their having the same mean. Combining both results renders as
$(J_{v}|J_{\mathrm{pa}(v)}=x_{\mathrm{pa(v)}})\preceq_{cx}(N_{v}^{\prime}|J_{\mathrm{pa}(v)}=x_{\mathrm{pa(v)}}),\quad\quad x_{\mathrm{pa(v)}}\in\{0,1\}.$ (30)
Consequently,
$((J_{\mathrm{pa}(v)}+J_{v})|J_{\mathrm{pa}(v)}=x_{\mathrm{pa}(v)})\preceq_{cx}((J_{\mathrm{pa}(v)}+N_{v}^{\prime})|J_{\mathrm{pa}(v)}=x_{\mathrm{pa}(v)}),\quad\quad x_{\mathrm{pa}(v)}\in\{0,1\}.$ (31)
Besides, a random variable $X$ is stochastically increasing convex (SICX) with respect to a random variable $\Theta$ if $\mathrm{E}[\varphi(X)|\Theta=\theta]$ is increasing convex in $\theta$ for every increasing convex function $\varphi$. Now, because the binomial distribution has the semigroup property in regard to its size parameter, the random variable $\alpha\circ X$, $\alpha\in[0,1]$, is SICX with respect to $X$. Example 8.A.7 of [Shaked and Shanthikumar, 2007] treats of the semigroup property of random variables and how it implies they are SICX. Next, let $N_{v}^{\ast}:=\alpha_{(\mathrm{pa}(v),v)}\circ\Theta_{\mathrm{pa}(v)}+L_{v}$, with $\Theta_{\mathrm{pa}(v)}$ acting as either $J_{\mathrm{pa}(v)}$, $N_{\mathrm{pa}(v)}^{\prime}$ or $N_{\mathrm{pa}(v)}$. Then, from the closure of SICX-ness on sums of conditionally independent random variables, as stated by Theorem 8.A.15 of [Shaked and Shanthikumar, 2007], $(\Theta_{\mathrm{pa}(v)}+\alpha_{(\mathrm{pa}(v),v)}\circ\Theta_{\mathrm{pa}(v)}+L_{v})$, i.e. $(\Theta_{\mathrm{pa}(v)}+N_{v}^{\ast})$, is SICX with respect to $\Theta_{\mathrm{pa}(v)}$, which, from Theorem 8.A.17 of [Shaked and Shanthikumar, 2007], the closure of SICX-ness on mixtures, implies $\left(\Theta_{\mathrm{pa}(v)}+N_{v}^{\ast}+\sum_{j\in\mathrm{dsc}(v)}N_{j}\right)$ is SICX with respect to $\Theta_{\mathrm{pa}(v)}$. Invoking again the closure of SICX-ness on sums of conditionally independent random variables, we deduce
$\left(\Theta_{u}+\sum_{i\in\mathrm{ch}(u)}\left(N_{i}^{\ast}+\sum_{j\in\mathrm{dsc}(i)}N_{j}\right)\right)\quad\text{is SICX with respect to}\;\Theta_{u},\,u\in\mathcal{V}.$ (32)
Theorem 8.A.14a of [Shaked and Shanthikumar, 2007] then implies, given (30), (32), and that both sides share the same expectation, for any $v\in\mathcal{V}\backslash\{r\}$,
$\left(\left.\left(J_{v}+\sum_{i\in\mathrm{ch}(v)}\left(N_{i}^{\prime}+\sum_{j\in\mathrm{dsc}(i)}N_{j}\right)\right)\right|J_{\mathrm{pa}(v)}=x_{\mathrm{pa(v)}}\right)\preceq_{cx}\left(\left.\left(N_{v}^{\prime}+\sum_{i\in\mathrm{ch}(v)}\left(N_{i}+\sum_{j\in\mathrm{dsc}(i)}N_{j}\right)\right)\right|J_{\mathrm{pa}(v)}=x_{\mathrm{pa(v)}}\right),$
for $x_{\mathrm{pa}(v)}\in\{0,1\}$. Let $\psi:=\max_{v\in\mathcal{V}}|\mathrm{path}(r,v)|$. Partition the vertices of the underlying tree in subsets $\{\Psi_{k},\;k\in\{0,\ldots,\psi\}\}$ of $\mathcal{V}$, depending of the length of their path to the root. That is $j\in\Psi_{i}$ if $|\mathrm{path}(r,j)|=i$. Then, in the same spirit as in (31) and given both models satisfies the global Markov property, we have, given (30),
$\left(\left.\left(\sum_{i=0}^{\psi}\sum_{j\in\Psi_{i}}J_{j}\right)\right|\bigcap_{i=0}^{\psi-1}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right)\preceq_{cx}\left(\left.\left(\sum_{i=0}^{\psi-1}\sum_{j\in\Psi_{i}}J_{j}+\sum_{k\in\Psi_{\psi}}N_{k}^{\prime}\right)\right|\bigcap_{i=0}^{\psi-1}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right),$ (34)
for $\bm{x}\in\{0,1\}^{d}$. By multiple applications of Theorem 3.A.12b of [Shaked and Shanthikumar, 2007], successively using every $J_{j}$, $j\in\Psi_{\psi-1}$, to act as the mixture random variable, (34) becomes
$\left(\left.\left(\sum_{i=0}^{\psi}\sum_{j\in\Psi_{i}}J_{j}\right)\right|\bigcap_{i=0}^{\psi-2}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right)$ $\preceq_{cx}\left(\left.\left(\sum_{i=0}^{\psi-1}\sum_{j\in\Psi_{i}}J_{j}+\sum_{k\in\Psi_{\psi}}N_{k}^{\prime}\right)\right|\bigcap_{i=0}^{\psi-2}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right)$ $\preceq_{cx}\left(\left.\left(\sum_{i=0}^{\psi-2}\sum_{j\in\Psi_{i}}J_{j}+\sum_{k\in\Psi_{\psi-1}}N_{k}^{\prime}+\sum_{l\in\Psi_{\psi}}N_{l}\right)\right|\bigcap_{i=0}^{\psi-2}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right),$ given (A.3). One then repeats the multiple applications of Theorem 3.A.12b, this time using every $J_{j}$, $j\in\Psi_{\psi-2}$, and obtains, for $\bm{x}\in\{0,1\}^{d}$,
$\left(\left.\left(\sum_{i=0}^{\psi}\sum_{j\in\Psi_{i}}J_{j}\right)\right|\bigcap_{i=0}^{\psi-3}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right)\preceq_{cx}\left(\left.\left(\sum_{i=0}^{\psi-2}\sum_{j\in\Psi_{i}}J_{j}+\sum_{k\in\Psi_{\psi-1}}N_{k}^{\prime}+\sum_{l\in\Psi_{\psi}}N_{l}\right)\right|\bigcap_{i=0}^{\psi-3}\bigcap_{j\in\Psi_{i}}\{J_{j}=x_{j}\}\right)$
[…]
again given (A.3). We repeat this argumentative $\psi-3$ other times, such that (A.3) unravels to
$\left(\left.\left(\sum_{i=0}^{\psi}\sum_{j\in\Psi_{i}}J_{j}\right)\right|\bigcap_{j\in\Psi_{0}}\{J_{j}=x_{j}\}\right)$ $\preceq_{cx}\left(\left.\left(\sum_{j\in\Psi_{0}}J_{j}+\sum_{k\in\Psi_{1}}N_{k}^{\prime}+\sum_{r=2}^{\psi}\sum_{l\in\Psi_{r}}N_{l}\right)\right|\bigcap_{j\in\Psi_{0}}\{J_{j}=x_{j}\}\right).$
Note that the sole vertex in subset $\Psi_{0}$ is the root. Applying one last time Theorem 3.A.12b and letting go of the $\Psi_{k}$ subsets notation, we obtain
$\left(\sum_{v\in\mathcal{V}}J_{v}\right)\leq_{cx}\left(J_{r}+\sum_{j\in\mathrm{ch(r)}}N_{j}^{\prime}+\sum_{v\in\mathcal{V}\backslash(\{r\}\cup\mathrm{ch}(r))}N_{v}\right).$ (36)
As exposed by (32), $\left(\Theta_{r}+\sum_{j\in\mathrm{ch(r)}}N_{j}^{\ast}+\sum_{v\in\mathcal{V}\backslash(\{r\}\cup\mathrm{ch}(r))}N_{v}\right)$ is SICX with respect to $\Theta_{r}$. Since $J_{r}\preceq_{cx}N_{r}$, for their respective distributions are Bernoulli and Poisson with identical means, applying Theorem 8.A.14a of [Shaked and Shanthikumar, 2007] yields
$\left(J_{r}+\sum_{j\in\mathrm{ch(r)}}N_{j}^{\prime}+\sum_{v\in\mathcal{V}\backslash(\{r\}\cup\mathrm{ch}(r))}N_{v}\right)\preceq_{cx}\left(\sum_{v\in\mathcal{V}}N_{v}\right).$ (37)
Joining (36) and (37), we deduce $\left(\sum_{v\in\mathcal{V}}J_{v}\right)\preceq_{cx}\left(\sum_{v\in\mathcal{V}}N_{v}\right)$. This is $K\preceq_{cx}M$. $\square$
Appendix B Natural parameters from joint probabilities
This appendix provides some intuition about the natural parameters $\bm{\eta}$ of a multivariate Bernoulli distribution by presenting the mapping to derive them from joint probabilities. Ensuing directly is the equivalence between (1) and (2), after a proper exponentiation. The following is Lemma 3.1 in [Dai et al., 2013]; we have restated it below following our notation.
Lemma 1 (Lemma 3.1 of [Dai et al., 2013]).
Assume $\bm{J}=(J_{v},\,v\in\mathcal{V})$ follows a multivariate Bernoulli distribution. Its natural parameters $\bm{\eta}=(\eta_{W},\,W\in\mathscr{V})$ are
$\eta_{W}=\ln\frac{\prod_{Z\subseteq W:|Z|\text{ is even}}\mathrm{Pr}\left(\bigcap_{v\in\mathcal{V}}\{J_{v}=\mathbbm{1}_{\{v\in Z\}}\}\right)}{\prod_{Z\subseteq W:|Z|\text{ is odd}}\mathrm{Pr}\left(\bigcap_{v\in\mathcal{V}}\{J_{v}=\mathbbm{1}_{\{v\in Z\}}\}\right)},\quad\text{for every }W\in\mathscr{V}.$
In particular, for the joint pmf in (2) to be normalized, one realizes that $A(\bm{\eta})=\eta_{\varnothing}=\ln\mathrm{Pr}(\bigcap_{u\in\mathcal{V}}\{J_{u}=0\})$. One still cannot derive $A(\bm{\eta})$ without summing over all combinations on $\{0,1\}^{d}$ if only (2) is given.
Appendix C Mean parameters from natural parameters
The objective of this appendix is not as much to provide a means to alternate between natural and mean parameterization than to than to show how untractable show how intricate such a mapping is, and thus highlight the unfixedness of the marginal distributions and pairwise correlations under the natural parameterization.
Let $\bm{J}=(J_{v},\,v\in\mathcal{V})$ be an Ising model defined on tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$. We derive a relation between the vector $\bm{\eta}$ of the natural parameterization (3) and the vectors $\bm{q}$ and $\bm{\alpha}$ of the mean parameterization (12), such that one may retrieve $\bm{q}$ and $\bm{\alpha}$ having $\bm{\eta}$. Let $\bm{x}_{-i}$ denote the vector $\bm{x}$ relieved of its $i$th element. We derive, for the marginal parameters, $i\in\mathcal{V}$,
$q_{i}=1-p_{J_{i}}(0)=1-\sum_{\bm{x}_{-i}\in\{0,1\}^{d-1}}p_{\bm{J}}(x_{1},\ldots,0,\ldots,x_{d}).$ (38)
We substitute (3) in (38) and obtain
$q_{i}$ $=1-\sum_{\bm{x}_{-i}\in\{0,1\}^{d-1}}\mathrm{e}^{\sum\limits_{v\in\mathcal{V}\backslash\{i\}}\eta_{(v,v)}x_{v}+\sum\limits_{\begin{subarray}{c}(u,v)\in\mathcal{E}\\ u,v\neq i\end{subarray}}\eta_{(u,v)}x_{u}x_{v}-A(\bm{\eta})}$
$=1-\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})}\sum_{\bm{x}_{-i}\in\{0,1\}^{d-1}}\mathrm{e}^{\sum\limits_{v\in\mathcal{V}\backslash\{i\}}\eta_{(v,v)}x_{v}+\sum\limits_{\begin{subarray}{c}(u,v)\in\mathcal{E}\\ u,v\neq i\end{subarray}}\eta_{(u,v)}x_{u}x_{v}-A^{*}_{i}(\bm{\eta})}$
$=1-\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})},$ (39)
where
$A^{*}_{i}(\bm{\eta})=\ln\sum_{\bm{y}\in\{0,1\}^{d-1}}\mathrm{e}^{\sum\limits_{v\in\mathcal{V}\backslash\{i\}}\eta_{(v,v)}y_{v}+\sum\limits_{\begin{subarray}{c}(u,v)\in\mathcal{E}\\ u,v\neq i\end{subarray}}\eta_{(u,v)}y_{u}y_{v}}.$
Then, for the correlation parameters, we have
$\alpha_{(i,j)}=\frac{p_{J_{i},J_{j}}(0,0)-(1-q_{i})(1-q_{j})}{\sqrt{q_{i}(1-q_{i})q_{j}(1-q_{j})}}.$ (40)
Substituting (39) in (40), we obtain
$\alpha_{(i,j)}$ $=\frac{p_{J_{i},J_{j}}(0,0)-(\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})})(\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})})}{\sqrt{\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})}(1-\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})})\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})}(1-\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})})}}$
$=\frac{\sum\limits_{\bm{x}_{-i,j}\in\{0,1\}^{d-2}}p_{\bm{J}}(x_{1},\ldots,0,\ldots,0,\ldots,x_{d})-(\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})})(\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})})}{\sqrt{\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})}(1-\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})})\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})}(1-\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})})}}$
$=\frac{\mathrm{e}^{A^{**}_{ij}(\bm{\eta})-A(\bm{\eta})}-(\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})})(\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})})}{\sqrt{\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})}(1-\mathrm{e}^{A^{*}_{i}(\bm{\eta})-A(\bm{\eta})})\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})}(1-\mathrm{e}^{A^{*}_{j}(\bm{\eta})-A(\bm{\eta})})}},$
where
$A^{**}_{ij}(\bm{\eta})=\ln\sum_{\bm{y}\in\{0,1\}^{d-1}}\mathrm{e}^{\sum\limits_{v\in\mathcal{V}\backslash\{i,j\}}\eta_{(v,v)}y_{v}+\sum\limits_{\begin{subarray}{c}(u,v)\in\mathcal{E}\\ u,v\not\in\{i,j\}\end{subarray}}\eta_{(u,v)}y_{u}y_{v}}.$
Let us note the arduous computations of $A^{*}_{ij}(\bm{\eta})$ and $A^{**}_{ij}(\bm{\eta})$ as they involve summation on every element of $\{0,1\}^{d}$. We remark that the marginal parameters and the dependence parameters are both functions of all components of $\bm{\eta}$, and even vary with respect to the shape on the tree, when relying on the canonical representation; the marginals are thus not fixed under this representation.